### v1

# CLIP-Guided Backdoor Defense through Entropy-Based Poisoned Dataset Separation

Abstract.
Deep Neural Networks (DNNs) are susceptible to backdoor attacks, where adversaries poison training data to implant backdoor into the victim model. Current backdoor defenses on poisoned data often suffer from high computational costs or low effectiveness against advanced attacks like clean-label and clean-image backdoors. To address them, we introduce CLIP-Guided backdoor Defense (CGD), an efficient and effective method that mitigates various backdoor attacks. CGD utilizes a publicly accessible CLIP model to identify inputs that are likely to be clean or poisoned. It then retrains the model with these inputs, using CLIP’s logits as a guidance to effectively neutralize the backdoor. Experiments on 4 datasets and 11 attack types demonstrate that CGD reduces attack success rates (ASRs) to below 1% while maintaining clean accuracy (CA) with a maximum drop of only 0.3%, outperforming existing defenses. Additionally, we show that clean-data-based defenses can be adapted to poisoned data using CGD. Also, CGD exhibits strong robustness, maintaining low ASRs even when employing a weaker CLIP model or when CLIP itself is compromised by a backdoor. These findings underscore CGD’s exceptional efficiency, effectiveness, and applicability for real-world backdoor defense scenarios. Code: https://github.com/binyxu/CGD.
Backdoor Defense, Contrastive Language-Image Pretraining
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/18/06submissionid: #845ccs: Computing methodologies Machine learningccs: Security and privacy Software and application security
1. Introduction
Deep Neural Networks (DNNs) are widely used in applications like facial recognition (An et al., 2023), autonomous driving (Han et al., 2022), and medical image diagnosis (Li et al., 2021c); however, backdoor attacks threaten their trustworthiness. By poisoning a small portion of the training data (Li et al., 2022a), adversaries can inject backdoors that cause models to make erroneous predictions when specific inputs are presented. Since the training data collection is usually time-consuming and expensive, it is common to use external data for training without security guarantees. The common practice makes backdoor attacks feasible in real-world applications, which highlights the importance of backdoor removal in poisoned datasets.
Existing backdoor defenses against poisoned data fall into two categories. (1) Clean model-based defenses (Huang et al., 2021; Gao et al., 2023; Chen et al., 2022) use self-/semi-supervised learning to train a clean model from the training dataset without relying on potentially compromised labels. These defenses identify backdoors by analyzing the behavior of this clean model. For example, DBD (Huang et al., 2021) employs self-supervised learning to train a clean encoder using only image data, whereas ASD (Gao et al., 2023) starts with a weak clean model and progressively improves it in a semi-supervised way. (2) Suspicious model-based defenses (Li et al., 2021b; Zheng et al., 2022) identify backdoors by analyzing the behavior of the potentially compromised model itself when processing with both benign and poisoned inputs. Techniques such as ABL (Li et al., 2021b) detect poisoned samples by noting their tendency for faster loss reduction during training. EP (Zheng et al., 2022), on the other hand, identifies “backdoor neurons” by observing that their pre-activation distributions differ significantly from those of benign neurons.
However, both categories all have their own limitations. (1) Clean model-based defenses are computationally intensive, often requiring hundreds to thousands of training epochs to train the clean model, given the high complexity of self/semi-supervised learning approach. Moreover, they struggle with advanced clean-label backdoors (Barni et al., 2019; Li et al., 2023), where the labels remain intact. In such scenarios, reducing the reliance on labels does not prevent the incorporation of backdoor information into the clean one. (2) Suspicious model-based defenses are less effective against clean-image backdoors (Jha et al., 2024; Xu et al., 2025), where the poisoned images are totally unchanged and share highly similar distributions from clean images. This similarity makes it challenging to detect anomalies based on the data distribution or the activations of the model.
To address these limitations, we proposed a novel defense strategy that leverages both clean and suspicious models from an entropy perspective. Our key insight is that a pretrained CLIP model, as a general zero-shot classifier, can serve as a weak but clean classifier across diverse tasks without additional training. By analyzing the cross-entropy score of predictions from CLIP and a suspicious model, we can effectively identify poisoned samples. Specifically, high entropy predictions from CLIP often indicate mislabeled samples, which is characteristic of label poisoning attacks. Conversely, in clean-label backdoors—where poisoned images are correctly labeled—the suspicious model tends to produce low-entropy predictions due to the strong association between the trigger and the target label.
Based on this observation, we introduce CLIP-Guided backdoor Defense (CGD), a two-step backdoor mitigation approach. (1) CLIP-Guided Meta Data Splitting: We generate a cross-entropy map (Fig. 1) by evaluating each sample with both CLIP and the suspicious model. Samples with high entropy predictions from CLIP and low entropy predictions from the suspicious model are flagged as potentially poisoned. This allows us to separate the dataset into clean and triggered subsets. (2) CLIP-Guided Backdoor Unlearning: Using the separated subsets, we retrain the model on clean data and perform unlearning on triggered data. During unlearning, CLIP guides the logits for the triggered samples, helping to remove the backdoor influence without degrading the model’s overall performance.
Figure 1. Entropy distribution plot for BadNets with CLIP as the weak but clean model. Label-poisoned samples (in red), including both normal and stealthy poisoned samples, are identifiable with high entropy in the clean model, while clean-label poisons (cl-poison in purple) are identifiable with low entropy in the suspicious model.
We conducted extensive experiments to validate the effectiveness of CGD. With a runtime under 3 minutes, CGD can impressively reduce attack success rates (ASRs) to 0.2%, 0.6%, 1.0%, and 0.1% on average across 11 different attack categories for CIFAR-10, CIFAR-100, GTSRB, and Tiny-ImageNet, respectively, outperforming all existing defenses. Notably, we innovatively apply clean-data-based defenses (Liu et al., 2018; Wu and Wang, 2021) to poisoned data by filtering a clean subset with CGD. Results show that defenses using poisoned data can even surpass those using 5% clean data. Remarkably, even with a very weak CLIP (7.5% accuracy), our defense can still achieve a 0.7% ASR and only a 0.7% drop in clean accuracy. Additionally, we demonstrate that even when CLIP contains its own backdoor, CGD can still remove the backdoor from the victim model without transferring CLIP’s backdoor. Our contributions are threefold:
  • Introduction of CLIP-Guided Defense (CGD): We proposed CGD, a novel backdoor defense mechanism that leverages CLIP as a weak but clean classifier, using entropy-based data separation and efficient unlearning to effectively remove backdoors from poisoned datasets.
  • Extensive Experimental Validation: Our method shows superior performance across multiple datasets and attack scenarios, reducing ASRs to under 1% and remaining robust even when using weak or backdoored CLIP models.
  • Introducing Clean-Data-Based Defenses on Poisoned Data: We introduce a novel insight that enables defenses traditionally dependent on clean data to operate on poisoned data by leveraging CGD to filter a clean subset. This concept opens up the possibility for clean-data defenses to be adapted for poisoned environments, offering a practical solution when clean data is unavailable.
2. Related Work
2.1. Multimodal Contrastive Learning (MCL)
Multimodal Contrastive Learning (MCL) bridges modalities by embedding large-scale data into a shared feature space. Specifically, in image-text MCL, it simultaneously learns visual and textual representations. CLIP (Radford et al., 2021), a seminal MCL model, achieves strong generalization by training on a 400M image-text dataset, treating paired images and texts as positives and all others as negatives. This robust cross-modal understanding has inspired advancements such as Uniclip (Lee et al., 2022), Cyclip (Goel et al., 2022) and DeCLIP (Li et al., 2022b). While most existing works on MCL focus on security concerns of CLIP, such as backdoor attacks (Carlini and Terzis, 2022; Liang et al., 2024) and defenses (Yang et al., 2024), our approach is novel. We leverage CLIP to enhance backdoor defense for other models, a direction not previously explored.
Figure 2. Pipeline for our CLIP-Guided Backdoor Defense (CGD).
2.2. Backdoor Attacks and Defenses
2.2.1. Backdoor Attacks.
Backdoor attacks are often implemented via data poisoning (Chen et al., 2017; Gu et al., 2019; Barni et al., 2019; Qi et al., 2022), where adversaries inject poison samples into the training dataset, causing the victim model to associate backdoor triggers with target classes. Beyond data poisoning, backdoor attacks can also manipulate the training process (Nguyen and Tran, 2020b, a; Wang et al., 2022), though these can often be adapted into data poisoning by introducing additional noise samples (Wu et al., 2024). Some attacks are designed specifically to evade defenses (Qi et al., 2022; Lin et al., 2020).
2.2.2. Clean-Data-Based Backdoor Defenses.
Existing defenses (Liu et al., 2018; Zhao et al., 2020; Wu and Wang, 2021; Zeng et al., 2021) often assume the availability of a small clean dataset (typically 5% of the training set) to mitigate backdoor attacks (Wu et al., 2024). Fine-Pruning (FP) (Liu et al., 2018) combines pruning and fine-tuning to remove backdoor-related neurons. Mode Connectivity Repair (MCR) (Zhao et al., 2020) navigates toward a clean model using clean data. Adversarial Neuron Pruning (ANP) (Wu and Wang, 2021) prunes neurons sensitive to triggers by adversarial training. Although effective, these methods require clean data, which may not always be available.
2.2.3. Poison-Data-Based Backdoor Defenses.
Poison-data-based defenses (Li et al., 2021b; Huang et al., 2021; Gao et al., 2023; Chen et al., 2022; Zheng et al., 2022) handle scenarios without clean data. (1) One approach focuses on representation learning. For instance, DBD (Huang et al., 2021) uses self-supervised learning to remove poisoned labels during pretraining, severing the link between the trigger and target label. D-BR (Chen et al., 2022) and ASD (Gao et al., 2023) further refine this by combining representation learning with semi-supervised approaches. However, these methods struggle against clean-label attacks, where triggers are embedded directly without modifying labels, allowing the victim model to still learn the trigger. (2) Another approach leverages the unique properties of triggered samples. Anti-Backdoor Learning (ABL) (Li et al., 2021b) reduces the influence of poisoned samples by down-weighting those with higher training losses. Entropy Pruning (EP) (Zheng et al., 2022) distinguishes backdoor neurons through unique bimodal pre-activation patterns. However, these methods are less effective against clean-image backdoors, where the images remain visually indistinguishable from clean ones.
Figure 3. Percentile rank of entropy for triggered images from intermediate models in clean-label backdoors (LC (Turner et al., 2019), SIG (Barni et al., 2019), CTRL (Li et al., 2023)). The 50% distribution range is marked with an area. Triggered images exhibit low entropy during training, enabling their identification.
3. Preliminary
3.0.1. Notations.
We consider a classification model $f(\cdot;\theta)$ parameterized by $\theta$. Let $p(y\mid x;\theta)$ denote the predicted probability assigned by the model $f(\cdot;\theta)$ to class $y$ given input $x$. The predicted label for $x$ is then defined as
$\hat{y}(x)=\arg\max_{y\in\mathcal{Y}}p(y\mid x;\theta),$
where $\mathcal{Y}=\{1,2,\ldots,C\}$ is the set of possible class labels. We define the trigger planting function, used to execute backdoor attacks, as $\mathcal{T}:\mathcal{X}\to\mathcal{X}$, and let $t\in\mathcal{Y}$ represent the attack’s designated target class. $\mathcal{P}$ denotes the distribution of clean, unaltered samples. The clean accuracy (CA) of the model is the probability of correct classification on clean data, $\mathbb{P}_{(x,y)\sim\mathcal{P}}[\hat{y}(x)=y]$. The attack success rate (ASR), $\mathbb{P}_{(x,y)\sim\mathcal{P}\mid y\neq t}[\hat{y}(\mathcal{T}(x))=t]$, measures the probability that a sample $x$ with an embedded trigger is misclassified as the target class $t$.
3.0.2. Threat model.
We adopt the poisoning-based threat model used in previous works (Chen et al., 2017; Gu et al., 2019; Turner et al., 2019), where attackers provide a poisoned training dataset containing a set of pre-created triggered samples. The defender neither knows the potential backdoor trigger pattern or even whether the dataset is poisoned. Moreover, the defender is believed to have access to a publicly available vision-languages model, such as CLIP (Radford et al., 2021). The goal of defenders is to obtain a well-performed model without suffering backdoor attacks.
4. Methodology
In this section, we introduce the pipeline of our CLIP-Guided Backdoor Defense (CGD) framework. As illustrated in Fig. 2, CGD comprises two main stages: the CLIP-Guided Meta Data Splitting stage and the CLIP-Guided Unlearning stage. Before deploying these stages, we assume the model owner has normally trained a suspicious model using a suspicious dataset that is potentially poisoned.
4.1. CLIP-Guided Meta Data Splitting
The novelty of our approach stems from leveraging CLIP, a pre-trained vision-language model, as a zero-shot classifier to detect backdoor samples in a poisoned dataset. Unlike traditional backdoor defenses that often rely on computationally intensive self-supervised or semi-supervised learning—approaches that struggle with certain backdoor types—our method exploits CLIP’s cross-modal understanding to identify subtle inconsistencies indicative of poisoning. This section outlines how we harness CLIP’s capabilities to address both poison-label and clean-label backdoors, offering a robust and efficient alternative to existing methods.
4.1.1. Poison-Label Backdoor Identification
Poison-label backdoors involve poisoned samples with incorrect or misleading labels, encompassing attacks such as classic backdoors (Gu et al., 2019; Chen et al., 2017), dynamic backdoors (Nguyen and Tran, 2020a, b; Wang et al., 2022), and clean-image backdoors (Jha et al., 2024; Xu et al., 2025). These attacks share a vulnerability: mislabeling that deviates from the true semantic content of the samples. We employ CLIP as a zero-shot classifier to detect such anomalies by computing entropy scores that reveal label inconsistencies.
For each sample $x_{i}$, CLIP generates a logit vector $\mathbf{z}_{i}^{\text{CLIP}}\in\mathbb{R}^{|L|}$ over a predefined label set $L$, without requiring additional training. We calculate the cross-entropy loss between CLIP’s predictions and the sample’s suspicious label $y_{i}$, yielding an entropy score $\mathcal{S}_{i}^{\text{CLIP}}$:
$\mathcal{S}_{i}^{\text{CLIP}}=-\log p(y_{i}\mid x_{i};\theta_{\text{CLIP}}),$
where $p(y_{i}\mid x_{i};\theta_{\text{CLIP}})$ is the probability CLIP assigns to the suspicious label $y_{i}$. A high $\mathcal{S}_{i}^{\text{CLIP}}$ suggests a mismatch between the image content and its label, flagging the sample as potentially poisoned. This approach leverages CLIP’s pre-trained alignment of visual and textual features, providing a novel, training-free mechanism for backdoor detection.
4.1.2. Clean-Label Backdoor Identification
Clean-label backdoors poison images without altering their original labels, making detection challenging. Here, triggered samples exhibit lower cross-entropy loss because the trigger acts as a dominant feature that the model learns to associate with the correct label during training. This heightened confidence reduces the loss, as illustrated in Fig. 3, where loss curves for attacks like LC (Turner et al., 2019), SIG (Barni et al., 2019), and CTRL (Li et al., 2023) drop and stabilize at low values. We exploit this property by thresholding the entropy scores from the suspicious model at epoch $T=5$, denoted as $\mathcal{S}_{i}^{\text{Model}}$:
$\mathcal{S}_{i}^{\text{Model}}=-\log p(y_{i}\mid x_{i};\theta_{\text{Model}}),$
where $p(y_{i}\mid x_{i};\theta_{\text{Model}})$ is the model’s predicted probability for the correct label. Samples with unusually low $\mathcal{S}_{i}^{\text{Model}}$ are flagged as potential clean-label backdoors, capitalizing on the model’s overconfidence induced by triggers.
4.1.3. Entropy Map Splitting
To ensure robustness across diverse attacks and datasets, we use percentile ranks of entropy scores, denoted by the function $\hat{\cdot}$, rather than raw values. For each sample $x_{i}$, we compute $\hat{\mathcal{S}}_{i}^{\text{CLIP}}$ and $\hat{\mathcal{S}}_{i}^{\text{Model}}$ from CLIP and the model, respectively. We then apply a threshold-based strategy with $\sigma_{1}$ and $\sigma_{2}$ to define the clean subset $D_{c}$ and triggered subset $D_{p}$:
$D_{c}=\left\{x_{i}\,\bigg{|}\,\hat{\mathcal{S}}_{i}^{\text{CLIP}}\leq 1-\sigma_{1}\ \text{and}\ \hat{\mathcal{S}}_{i}^{\text{Model}}\geq\sigma_{1}\right\},$
$D_{p}=\left\{x_{i}\,\bigg{|}\,\hat{\mathcal{S}}_{i}^{\text{CLIP}}>1-\sigma_{2}\ \text{or}\ \hat{\mathcal{S}}_{i}^{\text{Model}}<\sigma_{2}\right\}.$
The remaining samples form a mixed subset. To mitigate class imbalance in $D_{c}$, we oversample each class to match the original dataset’s label distribution.
4.2. CLIP-Guided Unlearning
Drawing on prior unlearning techniques (Li et al., 2021b; Chen et al., 2022; Gao et al., 2023), we fine-tune the backdoored model using a novel CLIP-guided approach. This stage integrates relearning, unlearning, and distillation to eliminate backdoor effects while preserving clean-data performance.
4.2.1. Relearning and Unlearning Losses
We retrain the model on $D_{c}$ with the standard cross-entropy loss $\mathcal{L}_{\text{re}}$ and apply a negative cross-entropy loss $\mathcal{L}_{\text{un}}$ on $D_{p}$:
$\mathcal{L}_{\text{re}}=-\sum_{x_{i}\in D_{c}}y_{i}\log p_{i},\quad\mathcal{L}_{\text{un}}=\sum_{x_{i}\in D_{p}}y_{i}\log(p_{i}+\epsilon),$
where $y_{i}$ is the true label, $p_{i}$ is the predicted probability, and $\epsilon$ prevents numerical instability.
4.2.2. CLIP-Guided Neural Distillation
To guide the model toward clean patterns, we introduce a distillation loss on $D_{p}$:
$\mathcal{L}_{\text{distill}}=\sum_{x_{i}\in D_{p}}\text{KL}\left(\mathbf{z}_{i}^{\text{CLIP}}\,\big{\|}\,\mathbf{z}_{i}^{\text{Model}}\right),$
[…]
Table 1. Poison data based defensive results on CIFAR-10 (CA and ASR) under poison rate of 5%.
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[100]Q[100]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48], cells = c, cell3-144 = bg = customred, cell3,4,8-116 = bg = customgreen, cell5-7,12-13,146 = bg = customred, cell3,5,7,8,12,138 = bg = customgreen, cell4,6,9-11,148 = bg = customred, cell3-5,7,8,1110 = bg = customgreen, cell6,9,10,12,13,1410 = bg = customred, cell3,4,8,1112 = bg = customgreen, cell5-7,9,10,12,13,1412 = bg = customred, cell3,6-1114 = bg = customgreen, cell4,5,12,13,1414 = bg = customred, cell3,4,6,7,8,1116 = bg = customgreen, cell5,9,10,12,13,1416 = bg = customred, cell3-1118 = bg = customgreen, cell12-1418 = bg = customred, cell3,4,6,7,10,1220 = bg = customgreen, cell5,8,9,11,13,1420 = bg = customred, cell3-1422 = bg = customgreen, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell111 = c=2, cell113 = c=2, cell115 = c=2, cell117 = c=2, cell119 = c=2, cell121 = c=2, cell21 = c=2, cell31 = r=2, cell51 = r=4, cell91 = r=3, cell121 = r=2, cell141 = c=2, cell151 = c=4, cell161 = c=4, hline1,17 = -0.10em, hline2 = 3,5,7,9,11,13,15,17,19,21l, hline2 = 4,6,8,10,12,14,16,18,20,22r, hline3,5,9,12,14-15 = -0.05em, Defense $\rightarrow$ & No Defense ABL (Li et al., 2021b) DBD (Huang et al., 2021) ASD (Gao et al., 2023) D-BR (Chen et al., 2022) EP (Zheng et al., 2022) ReBack (Ma et al., 2024) PIPD (Chen et al., 2024) MSPC (Pal et al., 2024) CGD (ours)
Attack $\downarrow$ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Classic Backdoor BadNets (Gu et al., 2019) 91.8 93.8 90.8 1.1 92.1 2.6 92.0 2.1 91.0 1.5 91.0 0.8 91.7 4.3 92.9 0.5 92.8 0.3 92.8 0.0
Blend (Chen et al., 2017) 93.7 99.8 86.2 4.4 93.0 100.0 93.0 5.3 85.1 0.0 91.6 96.1 91.7 2.4 93.1 5.3 92.7 0.7 93.2 0.0
Dynamic Backdoor WaNet (Nguyen and Tran, 2020b) 90.6 96.9 89.9 81.2 90.5 2.6 91.7 8.8 84.3 60.2 89.8 91.1 90.2 84.4 93.4 11.4 93.0 54.2 94.0 0.3
BPP (Wang et al., 2022) 91.4 99.2 89.1 99.8 92.4 99.9 92.5 99.4 88.5 85.5 89.8 4.6 90.1 1.8 93.1 0.9 90.5 2.8 94.0 0.3
IAB (Nguyen and Tran, 2020a) 89.7 94.9 89.3 83.0 91.3 0.0 92.3 19.8 85.3 84.8 90.7 1.6 87.9 1.7 92.0 4.0 92.5 5.3 93.3 0.7
SSBA (Li et al., 2021a) 93.0 97.3 88.6 3.9 92.5 2.4 93.3 7.1 83.1 3.0 92.0 10.5 85.1 6.6 90.6 17.2 90.9 21.5 92.9 0.0
Clean-Label Backdoor CTRL (Li et al., 2023) 93.6 95.9 88.2 2.4 92.1 57.8 91.3 89.3 90.5 98.3 92.3 1.1 91.4 96.2 90.1 12.6 91.8 77.3 93.6 0.1
SIG (Barni et al., 2019) 93.6 93.9 88.2 0.4 89.2 97.5 92.2 99.5 91.3 49.6 92.1 12.5 87.4 29.9 93.2 13.5 91.0 10.3 93.3 0.0
LC (Turner et al., 2019) 93.4 98.4 82.1 5.2 92.3 98.3 91.2 9.8 91.3 1.7 92.5 0.6 91.4 1.0 90.8 3.7 90.8 89.4 93.2 0.0
Clean-Image Backdoor FLIP (Jha et al., 2024) 89.9 99.2 84.8 99.6 92.3 1.6 86.9 62.2 83.9 22.1 89.9 98.5 90.0 39.7 91.4 66.9 91.6 17.2 92.5 0.5
GCB (Xu et al., 2025) 88.6 100.0 85.5 100.0 91.2 6.9 90.9 100.0 84.2 100.0 88.3 99.9 88.7 71.6 92.5 87.7 91.5 23.9 92.3 0.0
Average 91.7 97.2 87.5 43.7 91.7 42.7 91.6 45.7 87.2 46.1 90.9 37.9 89.6 30.7 92.1 20.3 91.7 27.5 93.2 0.2
CA Drop (smaller is better) $\downarrow$4.2 $\downarrow$0.0 $\downarrow$0.2 $\downarrow$4.6 $\downarrow$0.9 $\downarrow$2.1 $\downarrow$-0.4 $\downarrow$0.0 $\downarrow$-1.5
ASR Drop (larger is better) $\downarrow$53.5 $\downarrow$54.5 $\downarrow$51.5 $\downarrow$51.1 $\downarrow$59.3 $\downarrow$66.5 $\downarrow$76.9 $\downarrow$69.7 $\downarrow$97.0
4.2.3. Justification of Loss Terms
The three loss terms are meticulously designed to address distinct aspects of backdoor mitigation, with their necessity validated by ablation studies (see Section 5): Relearning Loss $\mathcal{L}_{\text{re}}$: This term reinforces the model’s ability to classify clean samples accurately, anchoring its performance on legitimate data. Without it, fine-tuning risks degrading generalization, as the model may overfit to the unlearning process. Unlearning Loss $\mathcal{L}_{\text{un}}$: By penalizing confident predictions on triggered samples, this term disrupts the trigger-label association critical to backdoor attacks. Its absence would leave the backdoor intact, as the model retains its poisoned behavior. Distillation Loss $\mathcal{L}_{\text{distill}}$: This term provides a positive learning signal, aligning the model’s predictions with CLIP’s clean, zero-shot classifications. Omitting it risks leaving the model without a coherent strategy for triggered samples, potentially reducing robustness.
Together, these terms form a synergistic framework: $\mathcal{L}_{\text{re}}$ preserves clean performance, $\mathcal{L}_{\text{un}}$ erases backdoor effects, and $\mathcal{L}_{\text{distill}}$ ensures a smooth transition to clean behavior, making the design principled rather than heuristic.
4.2.4. Final Loss Function
The total loss combines these terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{re}}+\lambda_{\text{un}}\mathcal{L}_{\text{un}}+\lambda_{\text{distill}}\mathcal{L}_{\text{distill}},$
where $\lambda_{\text{un}}$ and $\lambda_{\text{distill}}$ balance their contributions. Fine-tuning is limited to $K$ epochs, with early stopping if clean accuracy falls below $\tau$, ensuring performance preservation.
4.3. Extension: Enabling Clean-Data-Based Defenses on Poisoned Data
We extend our framework by using the split clean subset $D_{c}$ to enable clean-data-based defenses like Fine-Pruning (FP) (Liu et al., 2018), Mode Connectivity Repair (MCR) (Zhao et al., 2020), and Adversarial Neuron Pruning (ANP) (Wu and Wang, 2021) on poisoned datasets. Traditionally reliant on separate clean data, these methods gain practical utility through our approach, denoted as ANP-SC, MCR-SC, etc., enhancing their applicability and showcasing the versatility of our splitting strategy.
5. Evaluation
5.1. Experimental Setups
5.1.1. Attack Baselines and Setups.
We use four benchmark datasets: CIFAR-10 (Krizhevsky, 2009), CIFAR-100 (Krizhevsky, 2009), GTSRB (Stallkamp et al., 2012), and a subset of ImageNet (Russakovsky et al., 2015). PreActResNet18 (He et al., 2016) serves as the default model. We evaluate eleven state-of-the-art backdoor attacks across four categories: (1) Classic static-backdoor attacks: BadNets (Gu et al., 2019) and Blend (Chen et al., 2017); (2) Dynamic-backdoor attacks: SSBA (Li et al., 2021a), IAB (Nguyen and Tran, 2020a), WaNet (Nguyen and Tran, 2020b), and BPP (Wang et al., 2022); (3) Clean-label backdoor attacks: LC (Turner et al., 2019), SIG (Barni et al., 2019), and CTRL (Li et al., 2023); (4) Clean-image backdoor attacks: FLIP (Jha et al., 2024) and GCB (Xu et al., 2025). Attacks are implemented following guidelines from (Wu et al., 2022). We set the target label to 0 ($y_{t}=0$) and use a poison rate of 50% for clean-label attacks and 5% for the others.
5.1.2. Defenses.
We compare CGD with eight state-of-the-art established defenses: (1) Poisoned-data-based defenses: ABL (Li et al., 2021b), DBD (Huang et al., 2021), ASD (Gao et al., 2023), D-BR (Chen et al., 2022), EP (Zheng et al., 2022), ReBack (Ma et al., 2024), PIPD (Chen et al., 2024), and MSPC (Pal et al., 2024). D-BR, EP, ReBack, PIPD and MSPC use both the poisoned data and backdoored model, while ABL, DBD, and ASD only need poisoned data. These defenses can be integrated during normal training of the backdoored model on the poisoned dataset (2) Clean-data-based defenses: FP (Liu et al., 2018), MCR (Zhao et al., 2020), I-BAU (Zeng et al., 2021), and ANP (Wu and Wang, 2021). These methods rely on access to a clean subset, typically around 5% of the training data, as a baseline setting. We also evaluate these methods using clean data obtained from our CLIP-guided Splitting (SC) approach. For all methods, we use implementations from BackdoorBench whenever available; otherwise, we rely on their official implementations.
5.1.3. CGD Configuration.
For CGD, we use CLIP (Radford et al., 2021) with ViT-B-32, pretrained on Laion-2B (Schuhmann et al., 2022) via Open-CLIP. We set a uniform threshold $\sigma_{1}=0.1$ and $\sigma_{2}=0.2$ for triggered and clean data on all the experiments in evaluation except in specific ablation. Trade-off parameters are set as $\lambda_{\text{un}}=0.025$ and $\lambda_{\text{distill}}=0.0005$. We use $T=5$ as the intermediate epoch, and SGD with a learning rate of 0.01 and weight decay of $5\times 10^{-4}$ for fine-tuning.
5.1.4. Evaluation Metrics.
We evaluate defenses using clean accuracy (CA) and attack success rate (ASR). Specifically, ASR measures the fraction of non-target samples with triggers classified into the target label. An effective defense should maintain high CA while minimizing ASR, ensuring robustness against backdoor attacks without sacrificing model performance.
Table 2. Average training time (s) for poison-data-based defenses on CIFAR-10. Prepare refers to initial victim model training on poisoned data. Our method needs less than 3 minutes post victim model training.
{tblr}
width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[202]Q[200]Q[200]Q[200]Q[200], cells = c, hline1,11 = -0.10em, hline2 = -0.05em, Method & (Prepare) Data Split Defense Total
ABL (Li et al., 2021b) - 437 1,729 2,166
DBD (Huang et al., 2021) - 17,672 9,299 26,971
ASD (Gao et al., 2023) - 4,702 2,591 7,293
D-BR (Chen et al., 2022) 997 656 3,900 5,553
EP (Zheng et al., 2022) 997 - 53 1,050
ReBack (Ma et al., 2024) 997 112 92 1,201
PIPD (Chen et al., 2024) 997 1,325 438 2,760
MSPC (Pal et al., 2024) 997 1,844 352 3,193
CGD (ours) 997 60 98 1,154
Figure 4. Entropy distribution plot for various categories of attacks. Poisoned samples (red/purple) are easily separated.
5.2. Main Results
5.2.1. CLIP-Guided Defense (CGD).
To evaluate the effectiveness of CGD, we report the CA and ASR results for eight defense methods against eleven attacks on CIFAR-10 in Table 1. CGD consistently achieves low ASRs ($\leq 1\%$) while maintaining high CAs, with minimal drops ($\leq 0.3\%$) across all defenses. Remarkably, CGD can identify and unlearn poisoned samples, improving CA over the ”No Defense” baseline by 3.4% for WaNet and 3.6% for IAB. In contrast, self-supervised defenses like DBD, ASD, and D-BR fail to resist clean-label attacks such as CTRL and SIG, while property-based defenses like ABL and EP are vulnerable to clean-image backdoors like GCB and FLIP. Additionally, CGD is highly efficient, requiring less than 3 minutes to split data and apply defenses after training the backdoored model (Table 2). While EP (Zheng et al., 2022) takes a similar amount of time, it cannot separate clean data. Thus, CGD stands out as the most efficient method for isolating clean data from poisoned samples while maintaining a high defense success rate.
5.2.2. Entropy Map Visualization.
The entropy map distributions for the four defense categories, shown in Fig. 4 , highlight distinct patterns. (a) Classical backdoors (e.g., BadNets (Gu et al., 2019)): Poisoned samples cluster in the lower right, separable by entropy from suspicious models or CLIP. (b) Dynamic backdoors (e.g., IAB (Nguyen and Tran, 2020a)): Noise (green points) complicates detection but doesn’t inherently form backdoors. (c) Clean-label backdoors (e.g., SIG (Barni et al., 2019)): These resemble classical backdoors in suspicious model entropy, detectable by methods like ABL (Li et al., 2021b). (d) Clean-image backdoors (e.g., FLIP (Jha et al., 2024)): High entropy in suspicious models reduces the effectiveness of entropy- or loss-based defenses. Overall, each category’s unique entropy patterns support using category-average performance as a key metric in later experiments.
Table 3. Clean-data-based defenses. Methods with * use 5% clean data, while those with the “-SC” suffix use Split Clean data extracted from poisoned datasets using CGD.
{tblr}
width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[200]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87], cells = c, cell4,6,7,10,11,12,13,149 = bg = customgreen, cell3,59 = bg = customred, cell10,12,13,145 = bg = customgreen, cell3-7,115 = bg = customred, cell6,10,133 = bg = customgreen, cell3,4,5,7,11,12,143 = bg = customred, cell3,5,6,7,11,127 = bg = customgreen, cell4,10,13,147 = bg = customred, cell12,4,6,8 = c=20.174, cell82,4,6,8 = c=20.174, hline1,15 = -0.10em, hline2 = 2,4,6,8l, hline2 = 3,5,7,9r, hline3,8,10 = -0.05em, hline9 = 2,4,6,8l, hline9 = 3,5,7,9r, Defense $\rightarrow$ & FP* (Liu et al., 2018) MCR* (Zhao et al., 2020) I-BAU* (Zeng et al., 2021) ANP* (Wu and Wang, 2021)
Attack $\downarrow$ CA ASR CA ASR CA ASR CA ASR
Classic 92.1 21.7 92.6 51.1 86.4 16.5 85.3 21.4
Dynamic 92.8 30.9 72.6 47.9 90.5 30.3 87.2 1.1
Clean-Label 92.4 58.0 93.3 82.5 88.7 16.0 86.5 24.9
Clean-Image 92.1 17.7 92.9 38.5 89.4 6.6 84.9 0.3
Average 92.4 32.1 87.9 55.0 88.7 17.3 86.0 11.9
Defense $\rightarrow$ FP-SC MCR-SC I-BAU-SC ANP-SC
Attack $\downarrow$ CA ASR CA ASR CA ASR CA ASR
Classic 93.3 18.2 91.0 1.4 90.7 25.7 86.6 18.8
Dynamic 94.0 25.2 87.6 24.3 91.8 6.2 85.7 0.6
Clean-Label 93.2 57.2 90.8 5.7 90.8 17.8 87.5 9.7
Clean-Image 92.7 15.8 91.3 1.0 90.4 43.4 85.8 0.1
Average 93.3 29.1 90.2 8.1 90.9 23.3 86.4 7.3
5.2.3. Split Clean-Data Defense.
We incorporate a portion of selected clean data into clean-data-based defenses and observe a significant improvement in defense performance. Results comparing standard clean-data defenses with their counterparts using CLIP-guided data splitting are presented in Table 3. As shown, for most methods tested (FP (Liu et al., 2018), MCR (Zhao et al., 2020), ANP (Wu and Wang, 2021)), our Split Clean (SC) approach achieves superior outcomes, with both lower ASRs and higher CAs. This improvement is likely because our default hyperparameter setting allocates around 70% of the total data as clean, a significantly larger proportion than the typical clean-data assumption (around 5% of total data). Consequently, defenses such as ANP and MCR, which are based on the sensitivity of poisoned data, benefit from increased clean data in effectively mitigating backdoor effects. These results suggest that, for certain defenses, access to only 10% of poisoned data can yield defense results comparable to using 5% clean data, underscoring the effectiveness of the Split Clean approach.
5.2.4. Scalability.
Table 5.2.4 demonstrates the effectiveness of CGD across three additional datasets: CIFAR-100, GTSRB, and ImageNet. For each dataset, we evaluate all successful backdoor attacks (ASR $\geq$ 90%) and report the average results across attack categories. In all cases, CGD effectively eliminates backdoors from the models (with ASR reduced to $\leq$ 3%) while incurring only minimal clean accuracy (CA) reduction (average CA drop $\leq$ 0.3%). These results confirm the scalability and robustness of CGD across diverse datasets.
Table 4. CGD scalability on other datasets.
{tblr}
width = 0.8rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[250]Q[113]Q[113]Q[96]Q[119]Q[110]Q[110], cells = c, cell11 = r=2, cell12 = c=20.226, cell14 = c=20.215, cell16 = c=20.22, cell22 = c=20.226, cell24 = c=20.215, cell26 = c=20.22, cell91 = r=2, cell92 = c=20.226, cell94 = c=20.215, cell96 = c=20.22, cell102 = c=20.226, cell104 = c=20.215, cell106 = c=20.22, hline1,18 = -0.10em, hline3 = 2,4,6l, hline3 = 3,5,7r, hline11 = 2,4,6l, hline11 = 3,5,7r, hline4,9,12 = -0.05em, Dataset $\rightarrow$ & CIFAR-100  GTSRB  ImageNet
(No Defense) (No Defense) (No Defense)
Attack $\downarrow$ CA ASR CA ASR CA ASR
Classic 69.6 85.3 97.9 96.7 56.6 99.1
Dynamic 66.2 94.6 97.0 95.2 57.5 98.7
Clean-Label 70.6 38.3 98.4 85.5 51.6 48.1
Clean-Image 67.1 99.8 95.3 100.0 54.0 99.8
Average 68.4 79.5 97.1 94.4 54.9 86.4
Dataset $\rightarrow$ CIFAR-100  GTSRB  ImageNet
(CGD) (CGD) (CGD)
Attack $\downarrow$ CA ASR CA ASR CA ASR
Classic 69.6 0.0 97.8 0.0 56.5 0.0
Dynamic 70.4 0.3 97.8 1.1 57.7 0.0
Clean-Label 69.6 2.1 94.7 2.8 48.7 0.3
Clean-Image 68.2 0.0 96.8 0.0 55.6 0.0
Average 69.4 0.6 96.8 1.0 54.6 0.1
Average Drop -1.0 78.9 0.4 93.4 0.3 86.4
5.3. Ablation Study
5.3.1. Loss Term Analysis
In our CLIP-guided unlearning approach, we utilize three loss terms: unlearn loss ($\mathcal{L}_{\text{un}}$), relearn loss ($\mathcal{L}_{\text{re}}$), and distillation loss ($\mathcal{L}_{\text{distill}}$). To evaluate their roles, we conducted an ablation study by testing various combinations of these terms, with results shown in Table 5, measuring CA and ASR across multiple backdoor categories.
Single Loss Term. Using only $\mathcal{L}_{\text{un}}$ yields high CA (average: 88.1) but fails to reduce ASR (average: 93.8), notably against clean-label backdoors (ASR: 98.5), indicating insufficient backdoor mitigation without targeted guidance. Similarly, $\mathcal{L}_{\text{re}}$ alone maintains high CA (average: 92.9) yet struggles with high ASR (average: 86.3), as backdoor knowledge persists. Employing only $\mathcal{L}_{\text{distill}}$ reduces ASR (average: 33.3) but severely degrades CA (average: 34.1), making the model impractical.
Two-Term Combinations. Combining $\mathcal{L}_{\text{un}}$ and $\mathcal{L}_{\text{re}}$ improves defense (average ASR: 53.5), yet clean-label backdoors remain potent (ASR: 93.8), highlighting the need for additional guidance. Pairing $\mathcal{L}_{\text{re}}$ and $\mathcal{L}_{\text{distill}}$ reduces ASR significantly (average: 13.5), but dynamic backdoors retain a notable ASR (24.5), showing incomplete protection against sophisticated attacks.
Synergy of All Three Terms. Employing all three losses achieves optimal results, with an average CA of 93.1 and an ASR of 0.1 across all backdoor types. Here, $\mathcal{L}{\text{un}}$ targets backdoor unlearning, $\mathcal{L}{\text{distill}}$ aligns the model with correct logits via CLIP’s knowledge, and $\mathcal{L}_{\text{re}}$ preserves clean performance.
Figure 5. Threshold Study.
5.3.2. Hyperparameter Study.
The primary hyperparameters in our experiment are the clean data threshold $\sigma_{1}$ and the triggered data threshold $\sigma_{2}$, which define the splitting of the entropy map. Selecting suitable values for these thresholds is crucial: if set too low, triggered samples may be misclassified as clean, while overly high settings risk assigning clean samples to the triggered subset. Both cases reduce performance. To maintain $\sigma_{1}$ consistently higher than $\sigma_{2}$, we optimize using their ratio $\sigma_{1}/\sigma_{2}$. As illustrated in Fig. 5, ASR averaged over 11 attack methods highlights a robust range where ASR remains low. Even in worst-case scenarios, ASR is reduced below 30%. For default settings, we use $\sigma_{1}=0.2$ and $\sigma_{2}=0.1$, though these values can be flexibly adjusted within a reasonable range.
Table 5. Ablation study for different loss terms in CGD.
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[55]Q[55]Q[80]Q[60]Q[60]Q[65]Q[65]Q[74]Q[74]Q[71]Q[71]Q[60]Q[60], cells = c, cell11 = c=3, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell112 = c=2, vline4,12 = 1-90.05em, hline1,10 = -0.1em, hline3,9 = -0.05em, cell3-45,7,9,11,13 = bg=customred, cell55 = bg=customgreen, cell57,9,11,13 = bg=customred, cell65,7,9,11,13 = bg=customred, cell75,9,11 = bg=customgreen, cell77,13 = bg=customred, cell85,9,11,13 = bg=customgreen, cell87 = bg=customred, cell95,7,9,11,13 = bg=customgreen, Case Classic Dynamic Clean-label Clean-lmage Average
$\mathcal{L}_{\text{un}}$ $\mathcal{L}_{\text{re}}$ $\mathcal{L}_{\text{distill}}$ CA ASR CA ASR CA ASR CA ASR CA ASR
✓ 83.6 97.5 90.2 82.0 91.7 98.5 87.1 97.3 88.1 93.8
93.2 91.5 93.6 62.2 93.6 94.4 91.2 97.2 92.9 86.3
✓ 26.8 13.8 51.8 64.5 34.0 29.3 24.0 25.7 34.1 33.3
✓ ✓ 93.1 48.3 93.5 34.5 93.6 93.8 92.0 37.2 93.0 53.5
✓ ✓ 9.3 1.1 36.8 58.9 13.5 14.3 9.7 14.3 17.3 22.1
✓ ✓ 92.8 4.1 93.6 24.5 92.9 11.8 92.2 13.7 92.9 13.5
✓ ✓ ✓ 93.0 0.0 93.5 0.3 93.4 0.0 92.4 0.2 93.1 0.1
Table 6. CGD using other language-image pretraining models except CLIP.
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[260]Q[138]Q[138]Q[160]Q[160], cells = c, cell12 = c=2, cell14 = c=2, hline2 = 2,4l, hline2 = 3,5r, hline1,8 = -0.1em, hline3,7 = -0.05em, Model→ SigLIP ImageBind
Attack↓ CA ASR CA ASR
Classic 93.0 0.0 92.6 0.0
Dynamic 93.2 5.3 93.3 2.5
C-Label 92.8 0.7 92.7 0.7
C-Image 90.8 3.3 90.6 2.8
Average 92.5 2.3 92.3 1.5
5.3.3. Robustness to Different Language-Image Pretraining
Models
Although we mainly use CLIP (Radford et al., 2021) to illustrate CGD’s effectiveness, CGD can use other language-image pretraining models to replace CLIP. As shown in Table 6, we consider SigLIP (Zhai et al., 2023) and ImageBind (Girdhar et al., 2023) as two alternative models. The result suggests that both achieve low ASR ($\leq 5\%$) while maintaining good CA ($\geq 90\%$).
Figure 6. CGD’s defense performance on low-accuracy CLIP models. CGD successfully defends against all attacks (ASR $\leq 5\%$) even when CLIP’s zero-shot accuracy is as low as 10%.
5.4. Security Guarantee Analysis
Since our CGD is mainly based on the assumption of CLIP can act as an weak but clean model. As a result, there are two major security risks here: (A) If CLIP is a valid weak model for the specific task, and (B) If CLIP is a clean model. We evaluate these two assumptions one by one.
5.4.1. Weaker CLIP
We further investigate our defense’s robustness when CLIP’s accuracy is intentionally reduced. Given that current versions of CLIP are highly effective on common datasets, we manually introduce random noise to CLIP’s output, reducing its accuracy in a controlled manner. Specifically, we add Gumbel-distributed noise (Jang et al., 2016) to the output logits of CLIP, enabling us to adjust the level of inaccuracy by varying the scale parameter of the Gumbel distribution. As shown in Fig. 6, we present results illustrating the defense performance against the zero-shot classification accuracy of CLIP on CIFAR-100 and TinyImageNet. Our CGD approach successfully mitigates all tested attacks (ASR $\leq$ 5%) when CLIP’s zero-shot classification accuracy remains as low as 11.6% on CIFAR-100 and 7.5% on TinyImageNet. This is likely because even a weakened CLIP predicts images as similar classes close to the target label, causing the entropy-based calculations from CLIP to become less precise, though still somewhat informative. Additionally, while this noise does not impact backdoor removal capability, it does reduce clean accuracy in tandem with the drop in CLIP accuracy.
To further validate this assumption in real and diverse vision datasets, we carry out very extensive experiments on 17 different datasets, including large-scale datasets like SUN397 and full ImageNet. On these datasets, we tested three representative attacks (Blend, BPP, CTRL) but excluded clean-image backdoors due to either low ASR on large datasets (FLIP) or high computational cost (GCB). As shown in Table 7, CGD successfully defends all attacks except on SVHN for BPP. This failure is because CLIP’s zero-shot accuracy on SVHN is only 13.4%, making it thoroughly ineffective for a 10-class prediction dataset. Nonetheless, CGD still defends other SVHN attacks (e.g., Blend, CTRL) via the suspicious model’s entropy. Regardless of whether the dataset is large-scale or has low CLIP accuracy (e.g., GTSRB, DTD, and MNIST), CGD remains effective, showing its strong generalizability.
Table 7. CGD on 17 datasets against 3 backdoor attacks. “$\Delta$CA”: change in CA after defenses. CA drops $\leq$4.2%. CGD fails only when CLIP is thoroughly ineffective (e.g., on SVHN) on dynamic backdoor attack.
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[221]Q[92]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83], cells = c, cell11 = r=2, cell12 = r=3, cell13 = c=2,r=2, cell15 = c=6, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell43 = bg=customblue17, cell44 = bg=customblue0, cell53 = bg=customblue16, cell54 = bg=customblue17, cell63 = bg=customblue13, cell64 = bg=customblue0, cell73 = bg=customblue11, cell74 = bg=customblue7, cell83 = bg=customblue11, cell84 = bg=customblue5, cell93 = bg=customblue10, cell94 = bg=customblue0, cell103 = bg=customblue8, cell104 = bg=customblue9, cell113 = bg=customblue8, cell114 = bg=customblue9, cell123 = bg=customblue7, cell124 = bg=customblue7, cell133 = bg=customblue7, cell134 = bg=customblue5, cell143 = bg=customblue7, cell144 = bg=customblue5, cell153 = bg=customblue6, cell154 = bg=customblue2, cell163 = bg=customblue3, cell164 = bg=customblue1, cell173 = bg=customblue2, cell174 = bg=customblue1, cell183 = bg=customblue2, cell184 = bg=customblue5, cell193 = bg=customblue2, cell194 = bg=customblue1, cell203 = bg=customblue0, cell204 = bg=customblue0, cell4-206 = bg = customgreen, cell5-208 = bg = customgreen, cell48 = bg = customred, cell4-2010 = bg = customgreen, hline3 = 3,5,7,9,11l, hline3 = 4,6,8,10,12r, hline2 = 5-11, hline1,21 = -0.1em, hline4 = -0.05em, Task→ Data
Size
Classify
Accuracy
CGD (Ours)
Blend BPP CTRL
Dataset↓ CLIP RN50 $\Delta$CA ASR $\Delta$CA ASR $\Delta$CA ASR
SVHN 234M 13.4 95.7 0.5 0.1 -1.4 100 -4.2 8.6
Country211 20.9G 17.2 12.2 0.9 0.1 0.6 0.0 0.4 9.7
GTSRB 689M 32.6 98.5 -1.0 0.0 2.2 0.7 -3.8 0.2
FER2013 950M 41.4 62.1 -0.4 0.0 3.1 0.1 -0.4 0.0
DTD 1.2G 44.3 71.8 1.2 0.0 3.2 1.1 -2.2 0.7
MNIST 127M 48.2 99.6 -2.3 7.8 0.2 0.0 0.0 0.0
RenderedSST2 305M 58.6 53.8 2.8 0.0 4.8 0.5 -1.8 0.0
StanfordCars 3.7G 59.6 52.7 0.6 0.0 7.0 0.0 2.8 0.5
SUN397 73.6G 62.5 61.5 -0.4 0.0 2.0 0.0 0.2 5.5
ImageNet 146G 63.3 74.3 -0.3 0.0 0.1 0.0 -3.1 0.5
CIFAR100 338M 64.2 73.6 -0.4 0.0 4.9 0.1 -2.8 5.8
Flowers102 676M 66.5 89.5 -1.3 0.2 4.2 0.8 -1.3 0.4
Caltech101 324M 81.6 92.0 -2.9 3.6 5.5 3.0 -3.3 0.0
OxfordIIITPet 1.6G 87.3 92.9 0.8 0.3 8.3 12.4 1.9 0.0
Food101 5.3G 88.8 72.6 7.2 0.0 2.4 0.0 0.1 8.3
CIFAR10 340M 89.8 95.4 -0.5 0.0 2.6 0.3 0.0 0.1
STL10 5.4G 97.1 97.5 3.0 0.0 4.9 0.2 2.1 0.0
Table 8. CGD against backdoored CLIPs.
(i) (ii) (iii)
CA 92.8 92.3 92.3
ASR-O 0.0 0.0 0.1
ASR-C 0.0 0.6 0.1
5.4.2. Backdoored CLIP Analysis.
Previous research has shown that CLIP models can be vulnerable to backdoor attacks (Carlini and Terzis, 2022). In this study, we investigate whether a backdoored CLIP model could pass its backdoor to a victim model during our defense process. We explore three cases: (i) CLIP has a different backdoor trigger; (ii) CLIP has the same trigger but targets a different class; and (iii) CLIP has the same trigger and targets the same class. For each case, we use the BadNets trigger (Gu et al., 2019) (aligned with (Carlini and Terzis, 2022)) and measure two metrics: ASR-O (attack success rate for the suspicious dataset’s trigger) and ASR-C (attack success rate for CLIP’s trigger).
Our findings, shown in Table 8, reveal that even when CLIP is backdoored, it successfully removes backdoors from the victim model without transferring its own backdoor. This is mainly because the BadNets trigger, a simple and fixed pattern, can be detected by analyzing entropy in either CLIP or the victim model. As a result, the backdoor in CLIP does not compromise the defense, as the victim model’s entropy effectively identifies the trigger. However, in cases (ii) and (iii), we notice a small decrease in CA. This likely occurs because the poisoned outputs from CLIP introduce slight errors during the guidance process, mildly affecting CA.
5.5. Resistance to Potential Adaptive Attacks
5.5.1. Threat Model.
In the previous experiments, we assumed that attackers have no knowledge of our backdoor defense. In this section, we explore a more challenging scenario where attackers are aware of our defense strategy and adjust their poisoning strategies accordingly. However, they still cannot control the training process after injecting poisoned samples.
5.5.2. Methods.
Our CGD examines the poisoned dataset using both the suspicious model and a CLIP model. Attackers may evade detection by: (1) creating clean-label backdoors with high entropy in the suspicious model or (2) creating dynamic backdoors with low entropy in the CLIP model. To counter (1), we use an adaptive SIG-based trigger, Ada-SIG (Qi et al., 2022), enhanced by two techniques: Random Patch Masking (Rand-Patch), which divides the trigger into 16 patches and randomly masks 50%, and Random Opacity (Rand-Opacity), setting the trigger opacity to 50%. These methods make the trigger harder to detect. For (2), we apply the Feature Mixing Backdoor (FMB) attack (Lin et al., 2020), which mixes features from two classes to target a third class, confusing the CLIP model with multiple class features in one image.
5.5.3. Results.
Table 9 demonstrates that our CGD defense effectively mitigates both adaptive attacks. Against Adaptive-SIG, CGD reduces ASRs to $\leq 2.4\%$. As shown in Fig. 7, while some triggered samples evade detection during the splitting stage, they retain minimal trigger information, with strongly triggered samples all identified and unlearned. For the FMB attack, ASR is also reduced to $\leq 1.2\%$ across both one-to-one and all-to-one settings. This occurs because, although the CLIP model may confuse mixed-feature samples, it only misclassifies them within the two mixed classes, not the target (third) class. Consequently, the poisoned target label also yields a high cross-entropy score, making the attack ineffective against our defense.
Table 9. Potential adaptive attacks against our CGD.
{tblr}
width = 0.8rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[250]Q[250]Q[110]Q[110]Q[110]Q[110], cells = c, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell21 = c=2, cell31 = r=3, cell61 = r=2, hline1,8 = -0.1em, hline2 = 3,5l, hline2 = 4,6r, hline3,6 = -0.05em, Defense → No Defense CGD (Ours)
Adaptive Attacks ↓ CA ASR CA ASR
Ada-SIG Rand Patch 93.5 100.0 93.5 1.0
Rand Opacity 93.8 91.1 93.1 1.7
Both Rand 93.7 95.9 93.2 2.4
FMB One-to-One 92.0 85.0 92.0 0.3
All-to-One 91.5 77.7 91.2 1.2
6. Conclusion
We introduce CGD, a novel poison-data-based backdoor defense that leverages CLIP’s zero-shot capabilities to detect and mitigate attacks. Using an entropy-based method to separate poisoned data and guide model retraining, CGD reduces attack success rates below 1% across various datasets and attack types while maintaining high clean accuracy. Extensive experiments show CGD’s superior efficiency and resilience, even with weaker or backdoored CLIP models. CGD outperforms existing defenses and enhances clean-data-based methods, providing a practical and scalable solution for securing real-world deep learning applications.
Appendix A Detailed Mathematical Analysis of CLIP-Guided Backdoor Defense
This appendix presents a comprehensive mathematical analysis of the CLIP-Guided Backdoor Defense (CGD) framework, focusing on two core mechanisms: (1) detecting label-poisoned samples using CLIP’s zero-shot classification capabilities, and (2) identifying clean-label backdoors via the victim model’s cross-entropy scores. We derive probabilistic bounds using calibration properties and concentration inequalities, ensuring rigor and clarity.
A.1. Detection of Label-Poisoned Samples with CLIP
Consider a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, where a fraction $pr$ ($10<pr<$ 1) of samples are poisoned with incorrect labels. The clean subset is $C$ ($|C|=(1-pr)N$), and the poisoned subset is $P$ ($|P|=prN$). For each sample, CLIP computes a cross-entropy score $\mathcal{S}_{i}^{\text{CLIP}}=-\log p(y_{i}\mid x_{i})$, where $p(y_{i}\mid x_{i})$ is the predicted probability of the assigned label $y_{i}$. We rank samples by $\mathcal{S}_{i}^{\text{CLIP}}$ in descending order and flag the top $k=prN$ as potentially poisoned.
A.1.1. Calibration Assumptions and Score Behavior
CLIP’s effectiveness relies on its calibration, where predicted probabilities approximate true correctness likelihoods. The Expected Calibration Error (ECE) is defined as:
$\text{ECE}=\mathbb{E}_{p}\left[\left|\mathbb{P}(y=y_{\text{true}}\mid p(y\mid x)=p)-p\right|\right],$
assumed to be bounded by $\epsilon>0$, a small constant. With probability $1-\eta$ over the data distribution, uniform convergence ensures this holds for finite $N$, with a deviation term $\delta=O\left(\sqrt{\frac{\log(1/\eta)}{N}}\right)$.
For clean samples ($y_{i}=y_{\text{true},i}$), $p(y_{i}\mid x_{i})$ is typically high, yielding low $\mathcal{S}_{i}^{\text{CLIP}}$. For poisoned samples ($y_{i}\neq y_{\text{true},i}$), $p(y_{i}\mid x_{i})$ is low, resulting in high $\mathcal{S}_{i}^{\text{CLIP}}$. Define the cumulative distribution functions: - $F_{c}(s)=\mathbb{P}(\mathcal{S}_{i}^{\text{CLIP}}\leq s\mid x_{i}\in C)$, - $F_{p}(s)=\mathbb{P}(\mathcal{S}_{i}^{\text{CLIP}}\leq s\mid x_{i}\in P)$.
Calibration implies $F_{p}(s)\leq F_{c}(s)$, meaning poisoned scores stochastically dominate clean scores. We hypothesize a threshold $\theta$ where:
$\mathbb{P}(\mathcal{S}_{i}^{\text{CLIP}}\leq\theta\mid x_{i}\in C)=1-\alpha_{c},\quad\mathbb{P}(\mathcal{S}_{i}^{\text{CLIP}}>\theta\mid x_{i}\in P)=1-\alpha_{p},$
with $\alpha_{c},\alpha_{p}\leq\epsilon+\delta$.
A.1.2. Probabilistic Bound on Poisoned Sample Detection
Let $T$ denote the number of poisoned samples among the top $k=prN$ ranked by $\mathcal{S}_{i}^{\text{CLIP}}$. Define indicator variables: - $X_{i}=\mathbb{1}[\mathcal{S}_{i}^{\text{CLIP}}>\theta]$ for $x_{i}\in C$, with $\mathbb{P}(X_{i}=1)=p_{c}\leq\alpha_{c}$, - $Y_{j}=\mathbb{1}[\mathcal{S}_{j}^{\text{CLIP}}>\theta]$ for $x_{j}\in P$, with $\mathbb{P}(Y_{j}=1)=p_{p}\geq 1-\alpha_{p}$.
The total number of samples exceeding $\theta$ is $S=\sum_{i\in C}X_{i}+\sum_{j\in P}Y_{j}$, with expectations:
$\mathbb{E}\left[\sum_{i\in C}X_{i}\right]\leq(1-pr)N\alpha_{c},\quad\mathbb{E}\left[\sum_{j\in P}Y_{j}\right]\geq prN(1-\alpha_{p}).$
For the $k$-th threshold $\theta_{k}$ (where $k=prN$), the fraction of samples with $\mathcal{S}_{i}^{\text{CLIP}}>\theta_{k}$ satisfies:
$(1-pr)p_{c}+prp_{p}=pr.$
Solving, $p_{c}=\frac{pr(1-p_{p})}{1-pr}\leq\frac{pr\alpha_{p}}{1-pr}$. Thus, $T=\sum_{j\in P}\mathbb{1}[\mathcal{S}_{j}^{\text{CLIP}}>\theta_{k}]$, and:
$\mathbb{E}[T]=prNp_{p}\geq prN(1-\alpha_{p}).$
Using a Chernoff bound for $T$ (a sum of independent Bernoulli variables), for $t=(1-\gamma)prN$ where $\gamma>\alpha_{p}$:
$\mathbb{P}(T\leq t)\leq\exp\left(-\frac{(\mathbb{E}[T]-t)^{2}}{2\mathbb{E}[T]}\right).$
Substituting, $\mathbb{E}[T]-t=prN(p_{p}-(1-\gamma))\geq prN(\gamma-\alpha_{p})$, yielding:
$\mathbb{P}(T\leq(1-\gamma)prN)\leq\exp\left(-\frac{prN(\gamma-\alpha_{p})^{2}}{2(1-\alpha_{p})}\right).$
Thus, the success probability is:
$\mathbb{P}(T\geq(1-\gamma)prN)\geq 1-\exp\left(-\frac{prN(\gamma-\epsilon-\delta)^{2}}{2(1-\epsilon)}\right).$
This bound depends on the poisoning rate $pr$, sample size $N$, and CLIP’s calibration error $\epsilon$, confirming the method’s reliability.
A.2. Clean-Label Backdoor Detection via Victim Model
For clean-label backdoors, $prN$ samples contain a trigger but retain correct labels. The victim model, trained on this data, misclassifies triggered samples to a target class $t$. We compute $\mathcal{S}_{i}^{\text{Model}}=-\log p(y_{i}\mid x_{i})$, rank samples in ascending order, and flag the bottom $k=prN$ as poisoned.
A.2.1. Score Distribution and Backdoor Impact
For clean samples, $p(y_{i}\mid x_{i})$ is moderately high. For poisoned samples, the trigger induces $p(t\mid x_{i})\approx p_{t}$ (high confidence), reducing $p(y_{i}\mid x_{i})\leq 1-p_{t}$, thus lowering $\mathcal{S}_{i}^{\text{Model}}$. Assume:
$\mathbb{P}(\mathcal{S}_{i}^{\text{Model}}>\theta\mid x_{i}\in C)=1-\beta_{c},\quad\mathbb{P}(\mathcal{S}_{i}^{\text{Model}}\leq\theta\mid x_{i}\in P)=1-\beta_{p},$
where $\beta_{p}\leq 1-p_{t}$.
A.2.2. Bound on Detection Accuracy
Let $T$ be the number of poisoned samples in the bottom $k$. Define: - $X_{i}=\mathbb{1}[\mathcal{S}_{i}^{\text{Model}}\leq\theta]$ for $x_{i}\in C$, $\mathbb{P}(X_{i}=1)=p_{c}\leq\beta_{c}$, - $Y_{j}=\mathbb{1}[\mathcal{S}_{j}^{\text{Model}}\leq\theta]$ for $x_{j}\in P$, $\mathbb{P}(Y_{j}=1)=p_{p}\geq 1-\beta_{p}$.
For $\theta_{k}$ where $(1-pr)p_{c}+prp_{p}=pr$, we get $p_{c}\leq\frac{pr\beta_{p}}{1-pr}$, and:
$\mathbb{E}[T]=prNp_{p}\geq prN(1-\beta_{p}).$
Applying Chernoff bounds for $t=(1-\gamma)prN$:
$\mathbb{P}(T\leq(1-\gamma)prN)\leq\exp\left(-\frac{prN(\gamma-\beta_{p})^{2}}{2(1-\beta_{p})}\right),$
leading to:
$\mathbb{P}(T\geq(1-\gamma)prN)\geq 1-\exp\left(-\frac{prN(\gamma-(1-p_{t}))^{2}}{2p_{t}}\right).$
This bound highlights the dependence on $pr$, $N$, and trigger strength $p_{t}$, validating the approach’s effectiveness.
Appendix B Detailed Analysis of Resistance to Adaptive Attacks
In our main experiments, we operated under the assumption that attackers are unaware of our backdoor defense mechanisms. Here, we explore a more challenging scenario in which attackers have complete knowledge of our defense strategies and adapt their poisoning techniques accordingly.
B.1. Threat Model and Adaptive Attack Strategies
B.1.1. Threat Model.
Consistent with prior research (Chen et al., 2017; Gu et al., 2019; Turner et al., 2019), we assume attackers can access the entire training dataset and are informed about the defense mechanisms we employ. However, they lack control over the training process beyond injecting poisoned samples.
B.1.2. Adaptive Attack Strategies.
Our defense relies on examining both the suspicious model and a CLIP model, providing two potential avenues for attackers to attempt evasion:
  1. (1)
    Enhancing Clean-Label Backdoors’ Entropy in the Suspicious Model. Attackers aim to design triggers that produce higher entropy outputs in the suspicious model, making them less detectable. We incorporate the concept of Adaptive-Blend (Qi et al., 2022) with the SIG trigger (Barni et al., 2019), employing two enhancement techniques:
    • Random Patch Masking (Rand-Patch): The trigger is divided into 16 patches, with 50% randomly masked during training. This randomization reduces the trigger’s visibility and consistency.
    • Reduced Opacity (Rand-Opacity): The trigger’s opacity is set to 50%, making it less prominent in poisoned images.
    These methods aim to create stealthier triggers that are harder for our defense to isolate from benign samples.
  2. (2)
    Reducing Dynamic Backdoors’ Entropy in the CLIP Model. Attackers attempt to lower the cross-entropy scores in the CLIP model by using the Feature Mixing Backdoor (FMB) attack (Lin et al., 2020). This attack mixes features from different classes to create poisoned samples:
    • For instance, combining features from classes A and B to misclassify them as class C, while ensuring that unmixed features (e.g., A+A) are correctly classified.
    The FMB attack introduces ambiguity in the CLIP model’s predictions, potentially evading detection based on high cross-entropy scores.
Both adaptive attacks utilize a poison rate of 5%. For the FMB attack, an additional 50% of the samples are used for self-mixing noise to enhance stealthiness.
Figure 7. Entropy map for Adaptive SIG with both random patch and random opacity. Although a lot of samples are out of threshold line. Those samples are only weak poisoned samples with trivial trigger, while most of the strong triggered samples can be captions, leading to the success of our CGD.
Figure 8. Entropy map for Feature Mixing Backdoor (FMB) under the all-to-one setting: Data from the same class is considered its original label, while data from different classes is assigned the target label. Despite the combined features, all poisoned samples are easily identifiable under CLIP examination, leading to the success of our CGD.
(a) BadNets
(b) Blend
(c) SSBA
(d) WaNet
(e) BPP
(f) IAB
(g) CTRL
(h) SIG
(i) LC
(j) FLIP
(k) GCB
Figure 9. Entropy maps of various backdoor attack methods on the CIFAR-10 dataset. Each category of attack exhibits a distinct entropy distribution. Notably, SSBA, though classified as a dynamic backdoor attack, produces an entropy map resembling that of classic attacks. This similarity arises because SSBA does not introduce noise samples to prevent the victim network from learning trivial patterns. As a result, while the trigger appears dynamic, it behaves as a static pattern to the victim model.
B.2. Experimental Results and Analysis
Table 9 summarizes the effectiveness of our CGD defense against adaptive attacks. The results demonstrate strong resistance to both strategies.
  1. (1)
    Adaptive-SIG Methods: After applying our defense, the Attack Success Rate (ASR) drops significantly, remaining below 2.4%. Although some weakly poisoned samples with minimal trigger strength exceed our detection threshold, all strongly triggered samples—those posing the most significant threat—are effectively identified and unlearned (Fig. 7). This outcome indicates that while the attackers’ efforts increase the stealthiness of the triggers, our defense remains robust against the more impactful threats.
  2. (2)
    Feature Mixing Backdoor (FMB) Attack: The ASR after defense is consistently low, under 1.2% for both one-to-one and all-to-one scenarios. As shown in Fig. 8, the CLIP model may exhibit uncertainty when predicting correct labels for mixed-feature images, but it also assigns high cross-entropy scores to the attackers’ target labels. This dual uncertainty ensures that the poisoned samples are spotted by our defense mechanism, rendering the FMB attack ineffective.
B.2.1. Analysis.
The key to our defense’s effectiveness lies in its dual examination of model outputs and cross-entropy measures. For the Adaptive-SIG methods, the randomization and reduced opacity introduce variability that can occasionally slip past entropy-based detection but fail to protect stronger triggers, which are crucial for a successful attack. In the case of the FMB attack, the inherent ambiguity of mixed features undermines the attacker’s goal by preventing the model from confidently associating the poisoned samples with the target label without raising suspicion.
Appendix C More results on entropy map
The entropy maps in Fig. 9 reveal distinct patterns corresponding to four major backdoor attack categories: Classical Backdoors, Dynamic Backdoors, Clean-label Backdoors, and Clean-image Backdoors. For each attack type, we analyze the entropy distribution of both the weak model and a suspicious model trained on the poisoned dataset. These models provide insights into the separability of poisoned versus clean samples by their entropy values.
C.0.1. Classical backdoors
Classical backdoors, such as BadNets and Blend (Fig. 8(a) and 8(b)), exhibit a characteristic pattern where poisoned samples (marked in red) cluster in the lower right corner of the entropy map. This distribution suggests that poisoned samples have low entropy in the weak model while presenting higher entropy in the suspicious model. The sharp clustering in entropy space indicates that the classic poison samples can be separated effectively based on their entropy values. This separation enables entropy-based defenses, as these attacks leave a distinct and recognizable pattern that is challenging for the model to generalize across clean samples.
C.0.2. Dynamic backdoors
Dynamic backdoors, represented by methods such as IAB and BPP (Fig. 8(f) and 8(e)), involve the addition of noise or randomized triggers to poison samples. The entropy map for dynamic backdoors shows a more dispersed pattern, with green points representing noisy samples. These samples occupy a wider entropy range in both the weak and suspicious models. For example, IAB introduces noise to prevent the network from learning trivial patterns, leading to a distribution that is challenging to distinguish from clean data. Interestingly, SSBA (Fig. 8(c)), despite being categorized as a dynamic backdoor, displays a pattern more akin to classical backdoors. This discrepancy occurs because SSBA does not introduce noise samples, resulting in a static trigger behavior that the victim model interprets as a fixed pattern. Hence, SSBA behaves more like a classical backdoor from an entropy perspective.
C.0.3. Clean-label backdoors
Clean-label backdoors, such as SIG, CTRL, and LC (Fig. 8(h), 8(g), and 8(i)), exhibit a unique entropy pattern distinguishable primarily through the suspicious model’s entropy values. Clean-label attacks rely on naturally occurring features, blending triggers with benign data, leading to entropy values that overlap significantly with clean samples. The entropy maps for these attacks show that poisoned samples (marked in red) are less separable in the weak model. However, the suspicious model’s entropy values for these samples tend to deviate from clean ones, allowing for the application of entropy-based detection methods. Notably, entropy-based defenses like Anti-Backdoor Learning (ABL) (Li et al., 2021b) are effective against both clean-label and classical backdoors by leveraging these unique entropy distributions.
C.0.4. Clean-image Backdoors
Clean-image Backdoors, such as FLIP and GCB (Fig. 8(j) and 8(k)), present an entropy distribution that poses challenges for entropy-based defenses. These attacks use clean images with subtle perturbations, resulting in high entropy for both the weak and suspicious models. The distribution for clean-image backdoors is more uniform across both models, making it difficult to distinguish poisoned samples from clean ones based on entropy alone. This characteristic renders traditional entropy- and loss-based methods ineffective, as the high entropy across all samples blurs the distinction between clean and poisoned data. Consequently, novel defense approaches, potentially incorporating alternative metrics beyond entropy, may be required for robust detection of clean-image backdoors.
Appendix D Generalizability of Our Defense
To evaluate the robustness and generalizability of our defense mechanism, we conducted experiments under different poison rates and model architectures. The default setup for our evaluations uses a PreActResNet18 model with a 5% poison rate. However, we extend these tests to various scenarios to ensure that our defense is effective across a range of conditions.
D.1. Effectiveness Across Poison Rates
Table 10 illustrates the attack performance under different poison rates, specifically at 1% and 10%. In both cases, we observe that our defense (CGD) significantly reduces the Attack Success Rate (ASR) while maintaining a high Clean Accuracy (CA). For instance, at a 1% poison rate, the ASR for the BadNets attack is reduced from 73.8% without defense to 2.7% with our CGD defense. Similarly, at a 10% poison rate, CGD reduces the ASR for most attacks to below 1%, while the baseline ASR without defense remains above 90% for most attacks. This demonstrates that our defense is resilient even as the poison rate increases, effectively mitigating attacks and preserving model accuracy.
Table 10. Attack performance under different poison rates.
{tblr}
width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.10em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Poison Rate 1% 10%
Attack $\downarrow$ No Defense CGD (ours) No Defense CGD (ours)
CA ASR CA ASR CA ASR CA ASR
BadNets 93.2 73.8 92.6 2.7 91.8 93.8 92.3 0.0
Blend 93.8 94.1 93.5 0.0 93.7 99.8 93.5 0.0
WaNet 91.1 72.0 93.8 0.6 90.6 96.9 93.9 0.4
BPP 91.4 99.2 93.9 0.1 91.4 99.2 93.9 0.1
IAB 90.5 54.6 93.5 0.2 89.7 94.9 93.6 0.8
SSBA 93.4 99.7 93.3 0.4 93.0 97.3 93.0 0.0
CTRL 94.0 55.3 93.7 4.6 93.6 95.9 93.1 0.9
SIG 93.7 80.4 93.5 0.0 84.6 98.0 84.4 0.0
LC 93.5 68.3 93.2 2.8 84.4 99.8 84.4 0.0
FLIP 93.2 98.1 93.3 0.1 85.4 99.7 91.5 0.0
GCB 92.6 100.0 92.9 0.0 84.2 100.0 91.0 0.0
Average 92.8 81.4 93.4 1.0 89.3 97.8 91.3 0.2
Table 11. Attack performance under different architectures.
{tblr}
width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.08em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Architecture VGG16 MobileNet V2
Attack $\downarrow$ No Defense CGD (ours) No Defense CGD (ours)
CA ASR CA ASR CA ASR CA ASR
BadNets 89.5 95.0 90.0 0.0 82.1 90.5 81.5 1.7
Blend 90.7 98.1 91.0 0.0 82.3 96.9 81.9 1.7
WaNet 88.9 80.5 91.8 1.0 81.9 34.1 86.1 1.6
BPP 88.6 99.4 91.8 1.0 82.6 94.6 86.2 1.6
IAB 89.3 84.0 91.1 1.2 80.7 74.2 85.4 1.4
SSBA 89.9 80.7 90.4 0.6 80.3 68.3 80.5 1.8
CTRL 90.3 91.8 90.1 0.1 82.5 92.1 81.6 0.1
SIG 90.4 97.2 90.3 0.0 81.3 98.7 80.0 0.0
LC 90.3 87.8 90.0 0.0 82.6 96.4 81.6 0.1
FLIP 88.4 98.2 90.6 0.2 81.4 93.0 81.3 0.1
GCB 88.0 100.0 90.3 0.0 80.9 99.8 81.4 0.0
Average 89.5 92.1 90.7 0.4 81.7 85.3 82.5 0.9
D.2. Effectiveness Across Model Architectures
To further demonstrate the generalizability of our defense, we evaluated it on different model architectures, including VGG16 and MobileNet V2, as shown in Table 11. Our CGD defense consistently reduces the ASR across both architectures. For instance, the ASR for the BadNets attack is reduced from 95.0% to 0.0% on VGG16 and from 90.5% to 1.7% on MobileNet V2. Similar trends are observed across other attack methods, demonstrating the robustness of our defense across different neural network architectures.
In summary, our defense method (CGD) demonstrates strong generalizability, maintaining its effectiveness across a range of poison rates and architectures. These results underscore the robustness of our approach in mitigating attacks across diverse settings, making it suitable for a variety of practical applications.
Appendix E More Results in Ablation Study
We added an ablation here. Using only CLIP’s entropy mitigates most attacks (average ASR=13.8%) but struggles with clean-label attacks (46.4%). Model-only entropy excels there (ASR=1.3%) but falters elsewhere (average ASR=25.1%). Combining both yields the lowest average ASR (0.1%), as shown in Table 12.
Table 12. Ablation study for poisoned dataset splitting with entropy from different models.
{tblr}
width = 1.0rowsep = 0.5pt, colsep = 0.5pt, colspec = Q[150]Q[60]Q[60]Q[60]Q[60]Q[80]Q[80]Q[80]Q[80]Q[60]Q[60], cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell3-53 = bg=customgreen, cell3,55 = bg=customgreen, cell45 = bg=customred, cell4,57 = bg=customgreen, cell37 = bg=customred, cell3,59 = bg=customgreen, cell49 = bg=customred, cell3,511 = bg=customgreen, cell411 = bg=customred, vline2,10 = 1-50.05em, hline1,6 = -0.1em, hline3 = -0.05em, Category → Classic Dynamic Clean-label Clean-image Average
Method ↓ CA ASR CA ASR CA ASR CA ASR CA ASR
CLIP only 92.8 0.6 93.4 4.7 92.8 46.4 92.2 3.6 92.8 13.8
Victim only 92.8 1.0 92.7 57.7 93.3 1.3 90.4 38.7 92.3 25.1
Both 93.0 0.0 93.5 0.3 93.4 0.0 92.4 0.2 93.1 0.1
Appendix F Use CLIP Variants in Domain-Specific Tasks
CGD’s reliance on a CLIP-like model is practical because defenders can use domain-specific CLIP variants for specific tasks. Across 17 datasets tested, CGD succeeded in 16, showcasing its broad applicability. For the sole exception, SVHN (vanilla CLIP accuracy: 13.4%, near-random guess for 10-class dataset), switching to LAION-2B CLIP variant reduced the ASR to 2.6% under the BPP attack.
In other domain-specific environments like medical imaging, CLIP models such as PubMed (Eslami et al., 2023) (62.7% accuracy on PCAM) achieved an ASR of 6.1%, as shown in Table 13. These results demonstrate that with advanced domain-specific CLIP variants (e.g. PubMedCLIP (Eslami et al., 2023) in Medicine, BioCLIP (Stevens et al., 2024) in Biology, GeomCLIP (Xiao et al., 2024) in Chemistry), defenders can select high-performing CLIP variants on the poisoned dataset to ensure CGD’s effectiveness. CLIP’s public availability and ongoing advancements in domain-specific models further reduce feasibility concerns.
Table 13. Performance of different CLIP pre-training weights on SVHN and PCAM.
{tblr}
width = 0.8rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[238]Q[254]Q[192]Q[113]Q[113], cells = c, cell12 = r=2, cell13 = r=2, cell14 = c=2, cell31 = r=3, cell61 = r=3, hline1,9 = -0.1em, hline3,6 = -0.05em, Task→ CLIP Weight CLIP Acc. BPP Attack
Dataset↓ CA  ASR
SVHN
(10 classes) Vanilla 13.4 94.3 100
LAION-400M (Schuhmann et al., 2022) 27.7 94.8 15.2
LAION-2B (Schuhmann et al., 2022) 38.8 95.6 2.6
PCAM
(2 classes) Vanilla 52.3 82.1 69.5
BioMed (Zhang et al., 2025) 60.1 83.8 15.8
PubMed (Eslami et al., 2023) 62.7 85.4 6.1
Table 14. Clean-data-based defenses (5%) vs poisoned-data-based defenses (100%). In this case, poisoned-data-based defenses can get better defense performance than clean-data-based defenses with the help of “SC” (Split Clean data).
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,12-1412 = bg = customgreen, cell5,7,10,11,1512 = bg = customred, cell4,6-1514 = bg = customgreen, cell514 = bg = customred, cell4,6-9,11-13,1516 = bg = customgreen, cell5,10,1416 = bg = customred, cell4-13,1518 = bg = customgreen, cell1418 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 100% Poisoned Training Data
Defense → FP (Liu et al., 2018) ANP (Wu and Wang, 2021) I-BAU (Zeng et al., 2021) MCR (Zhao et al., 2020) FP-SC ANP-SC I-BAU-SC MCR-SC
Attack ↓ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Classic
Backdoor BadNets 92.3 2.1 85.6 0.0 86.2 0.8 92.1 2.4 93.3 1.0 85.4 0.0 90.1 1.5 91.2 1.0
Blend 91.9 41.4 85.0 42.7 86.5 32.3 93.1 99.8 93.3 35.3 87.9 37.6 91.3 49.9 90.7 1.9
Dynamic
Backdoor WaNet 93.2 5.4 82.6 0.6 90.3 3.8 93.4 12.9 94.5 2.1 85.5 0.1 91.0 5.1 92.1 1.1
BPP 93.2 66.9 82.7 0.6 90.9 91.6 93.4 2.0 94.3 89.0 87.3 0.5 92.7 13.1 91.7 2.2
IAB 92.9 36.8 85.5 1.3 90.8 24.2 93.0 53.6 94.1 2.1 90.6 0.3 92.3 2.4 91.7 2.5
SSBA 92.0 14.5 87.3 0.1 90.0 1.5 92.4 64.2 92.9 7.7 86.1 0.0 91.2 4.3 90.5 1.0
Clean-Label
Backdoor CTRL 92.5 98.5 93.0 3.8 89.5 25.4 93.5 99.3 93.1 78.5 86.1 1.9 90.2 41.2 90.6 2.5
SIG 92.7 40.9 86.3 22.5 87.7 17.5 93.4 93.9 93.2 73.9 86.3 12.1 90.1 10.6 90.3 13.2
LC 92.1 34.7 87.6 51.1 88.9 5.3 93.6 100.0 93.1 19.3 85.5 16.7 92.1 1.6 90.4 1.4
Clean-Image
Backdoor FLIP 91.9 0.1 91.7 0.1 87.9 0.3 91.9 0.3 93.1 12.7 88.0 0.0 90.1 0.3 90.6 0.3
GCB 92.3 35.4 81.2 0.0 90.9 12.8 91.2 90.0 92.4 19.0 81.4 0.0 90.7 86.4 90.3 92.4
Average 92.5 34.2 86.2 11.2 89.1 19.6 92.8 56.2 93.4 31.0 86.4 6.3 91.1 19.7 90.9 10.9
Table 15. Clean-data-based defenses (5%) vs poisoned-data-based defenses (10%). In this case, poisoned-data-based defenses can get comparable defense performance to clean-data-based defenses with the help of “SC” (Split Clean data).
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,1312 = bg = customgreen, cell5,7,10-12,14,1512 = bg = customred, cell4,6-11,13-1514 = bg = customgreen, cell5,1214 = bg = customred, cell4,7-13,1516 = bg = customgreen, cell5,6,1416 = bg = customred, cell6-818 = bg = customgreen, cell4,5,9-1518 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 10% Poisoned Training Data
Defense → FP (Liu et al., 2018) ANP (Wu and Wang, 2021) I-BAU (Zeng et al., 2021) MCR (Zhao et al., 2020) FP-SC ANP-SC I-BAU-SC MCR-SC
Attack ↓ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Classic
Backdoor BadNets 92.3 2.1 85.6 0.0 86.2 0.8 92.1 2.4 92.5 2.7 85.0 0.0 89.5 1.2 92.8 34.1
Blend 91.9 41.4 85.0 42.7 86.5 32.3 93.1 99.8 92.5 37.6 88.4 42.5 89.0 59.4 93.5 99.3
Dynamic
Backdoor WaNet 93.2 5.4 82.6 0.6 90.3 3.8 93.4 12.9 89.6 2.1 82.2 0.1 91.8 22.1 93.1 0.6
BPP 93.2 66.9 82.7 0.6 90.9 91.6 93.4 2.0 93.2 90.4 88.1 0.8 91.7 9.0 92.9 2.8
IAB 92.9 36.8 85.5 1.3 90.8 24.2 93.0 53.6 93.3 1.9 83.7 0.0 91.8 2.9 92.9 1.2
SSBA 92.0 14.5 87.3 0.1 90.0 1.5 92.4 64.2 88.6 16.1 86.0 0.0 90.2 10.9 92.5 51.5
Clean-Label
Backdoor CTRL 92.5 98.5 93.0 3.8 89.5 25.4 93.5 99.3 91.1 95.3 93.2 2.3 90.5 6.5 93.6 99.6
SIG 92.7 40.9 86.3 22.5 87.7 17.5 93.4 93.9 92.6 55.7 88.1 7.3 88.6 7.6 93.5 97.8
LC 92.1 34.7 87.6 51.1 88.9 5.3 93.6 100.0 92.1 47.8 89.3 56.6 88.5 7.1 92.9 100.0
Clean-Image
Backdoor FLIP 91.9 0.1 91.7 0.1 87.9 0.3 91.9 0.3 92.1 1.9 90.5 0.0 88.2 3.4 92.5 76.7
GCB 92.3 35.4 81.2 0.0 90.9 12.8 91.2 90.0 92.1 59.5 84.5 3.2 89.3 87.6 93.2 89.1
Average 92.5 34.2 86.2 11.2 89.1 19.6 92.8 56.2 91.8 37.4 87.2 10.3 89.9 19.8 93.0 59.3
Appendix G More Results on Defenses Based on Clean Data
As mentioned in the introduction, we introduce a novel insight that enables defenses traditionally dependent on clean data to operate effectively on poisoned data by leveraging CGD to filter a clean subset. This concept opens up the possibility for adapting clean-data defenses for poisoned environments, offering a practical solution when clean data is unavailable. In this section, we evaluate clean-data based defenses under three conditions: 5% clean training dataset, whole (100%) poisoned training dataset, and 10% poisoned training dataset.
Table 14 presents the results of clean-data based defenses with 5% clean training data available. In this setting, the total number of successful defenses (defined as defenses achieving an Attack Success Rate (ASR) of 20% or lower, highlighted in green) is 23.
Table 14 shows the results for clean-data based defenses when a clean subset is split from the whole (100%) poisoned training dataset. With access to the full poisoned dataset, the defense performance improves significantly, with the number of successful defenses increasing to 35, marking a 50% improvement over the 5% clean data setting. This indicates that having access to the full poisoned dataset allows the defenses to operate more effectively, even in the absence of originally clean data.
Table 15 provides results for clean-data based defenses with a 10% poisoned training dataset. Even in this setting, where only a small portion of the dataset is clean, the total number of successful defenses increases from 23 (in the 5% clean data setting) to 25. This demonstrates that a 10% poisoned dataset can yield comparable performance to that obtained with 5% clean data, suggesting that even minimal clean data availability can significantly bolster defense effectiveness.
In summary, our findings suggest that accessing the full (100%) poisoned dataset enables all tested defense methods to achieve better performance. The total number of successful defenses (ASR $\leq$ 20%) increased from 23 to 35, representing a 50% improvement over the original method that used only 5% clean data. Even with a 10% poisoned dataset, the number of successful defenses increased from 23 to 25, indicating that 10% poisoned data achieves a comparable performance to 5% clean data. This demonstrates the potential of clean-data-based defenses to perform effectively in poisoned environments, even with limited clean data availability.

### v2

@@
 Backdoor Defense, Contrastive Language-Image Pretraining
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/18/06submissionid: #845ccs: Computing methodologies Machine learningccs: Security and privacy Software and application security
+journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754860isbn: 979-8-4007-2035-2/2025/10submissionid: #845ccs: Computing methodologies Machine learningccs: Security and privacy Software and application security
 1. Introduction
@@
 Table 1. Poison data based defensive results on CIFAR-10 (CA and ASR) under poison rate of 5%.
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[100]Q[100]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48], cells = c, cell3-144 = bg = customred, cell3,4,8-116 = bg = customgreen, cell5-7,12-13,146 = bg = customred, cell3,5,7,8,12,138 = bg = customgreen, cell4,6,9-11,148 = bg = customred, cell3-5,7,8,1110 = bg = customgreen, cell6,9,10,12,13,1410 = bg = customred, cell3,4,8,1112 = bg = customgreen, cell5-7,9,10,12,13,1412 = bg = customred, cell3,6-1114 = bg = customgreen, cell4,5,12,13,1414 = bg = customred, cell3,4,6,7,8,1116 = bg = customgreen, cell5,9,10,12,13,1416 = bg = customred, cell3-1118 = bg = customgreen, cell12-1418 = bg = customred, cell3,4,6,7,10,1220 = bg = customgreen, cell5,8,9,11,13,1420 = bg = customred, cell3-1422 = bg = customgreen, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell111 = c=2, cell113 = c=2, cell115 = c=2, cell117 = c=2, cell119 = c=2, cell121 = c=2, cell21 = c=2, cell31 = r=2, cell51 = r=4, cell91 = r=3, cell121 = r=2, cell141 = c=2, cell151 = c=4, cell161 = c=4, hline1,17 = -0.10em, hline2 = 3,5,7,9,11,13,15,17,19,21l, hline2 = 4,6,8,10,12,14,16,18,20,22r, hline3,5,9,12,14-15 = -0.05em, Defense $\rightarrow$ & No Defense ABL (Li et al., 2021b) DBD (Huang et al., 2021) ASD (Gao et al., 2023) D-BR (Chen et al., 2022) EP (Zheng et al., 2022) ReBack (Ma et al., 2024) PIPD (Chen et al., 2024) MSPC (Pal et al., 2024) CGD (ours)
+Defense $\rightarrow$ No Defense ABL (Li et al., 2021b) DBD (Huang et al., 2021) ASD (Gao et al., 2023) D-BR (Chen et al., 2022) EP (Zheng et al., 2022) ReBack (Ma et al., 2024) PIPD (Chen et al., 2024) MSPC (Pal et al., 2024) CGD (ours)
 Attack $\downarrow$ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
@@
 Our method needs less than 3 minutes post victim model training.
-{tblr}
-width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[202]Q[200]Q[200]Q[200]Q[200], cells = c, hline1,11 = -0.10em, hline2 = -0.05em, Method & (Prepare) Data Split Defense Total
+Method (Prepare) Data Split Defense Total
 ABL (Li et al., 2021b) - 437 1,729 2,166
@@
 5.2.1. CLIP-Guided Defense (CGD).
-To evaluate the effectiveness of CGD, we report the CA and ASR results for eight defense methods against eleven attacks on CIFAR-10 in Table 1.
-CGD consistently achieves low ASRs ($\leq 1\%$) while maintaining high CAs, with minimal drops ($\leq 0.3\%$) across all defenses.
-Remarkably, CGD can identify and unlearn poisoned samples, improving CA over the ”No Defense” baseline by 3.4% for WaNet and 3.6% for IAB.
-In contrast, self-supervised defenses like DBD, ASD, and D-BR fail to resist clean-label attacks such as CTRL and SIG, while property-based defenses like ABL and EP are vulnerable to clean-image backdoors like GCB and FLIP.
-Additionally, CGD is highly efficient, requiring less than 3 minutes to split data and apply defenses after training the backdoored model (Table 2).
-While EP (Zheng et al., 2022) takes a similar amount of time, it cannot separate clean data.
-Thus, CGD stands out as the most efficient method for isolating clean data from poisoned samples while maintaining a high defense success rate.
+[…]
 5.2.2. Entropy Map Visualization.
@@
 Table 3. Clean-data-based defenses. Methods with * use 5% clean data, while those with the “-SC” suffix use Split Clean data extracted from poisoned datasets using CGD.
-{tblr}
-width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[200]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87], cells = c, cell4,6,7,10,11,12,13,149 = bg = customgreen, cell3,59 = bg = customred, cell10,12,13,145 = bg = customgreen, cell3-7,115 = bg = customred, cell6,10,133 = bg = customgreen, cell3,4,5,7,11,12,143 = bg = customred, cell3,5,6,7,11,127 = bg = customgreen, cell4,10,13,147 = bg = customred, cell12,4,6,8 = c=20.174, cell82,4,6,8 = c=20.174, hline1,15 = -0.10em, hline2 = 2,4,6,8l, hline2 = 3,5,7,9r, hline3,8,10 = -0.05em, hline9 = 2,4,6,8l, hline9 = 3,5,7,9r, Defense $\rightarrow$ & FP* (Liu et al., 2018) MCR* (Zhao et al., 2020) I-BAU* (Zeng et al., 2021) ANP* (Wu and Wang, 2021)
+Defense $\rightarrow$ FP* (Liu et al., 2018) MCR* (Zhao et al., 2020) I-BAU* (Zeng et al., 2021) ANP* (Wu and Wang, 2021)
 Attack $\downarrow$ CA ASR CA ASR CA ASR CA ASR
@@
 5.2.4. Scalability.
-Table 5.2.4 demonstrates the effectiveness of CGD across three additional datasets: CIFAR-100, GTSRB, and ImageNet.
+Table 4 demonstrates the effectiveness of CGD across three additional datasets: CIFAR-100, GTSRB, and ImageNet.
 For each dataset, we evaluate all successful backdoor attacks (ASR $\geq$ 90%) and report the average results across attack categories.
@@
 Table 4. CGD scalability on other datasets.
-{tblr}
-width = 0.8rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[250]Q[113]Q[113]Q[96]Q[119]Q[110]Q[110], cells = c, cell11 = r=2, cell12 = c=20.226, cell14 = c=20.215, cell16 = c=20.22, cell22 = c=20.226, cell24 = c=20.215, cell26 = c=20.22, cell91 = r=2, cell92 = c=20.226, cell94 = c=20.215, cell96 = c=20.22, cell102 = c=20.226, cell104 = c=20.215, cell106 = c=20.22, hline1,18 = -0.10em, hline3 = 2,4,6l, hline3 = 3,5,7r, hline11 = 2,4,6l, hline11 = 3,5,7r, hline4,9,12 = -0.05em, Dataset $\rightarrow$ & CIFAR-100  GTSRB  ImageNet
+Dataset $\rightarrow$ CIFAR-100 GTSRB ImageNet
 (No Defense) (No Defense) (No Defense)
@@
 Table 5. Ablation study for different loss terms in CGD.
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[55]Q[55]Q[80]Q[60]Q[60]Q[65]Q[65]Q[74]Q[74]Q[71]Q[71]Q[60]Q[60], cells = c, cell11 = c=3, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell112 = c=2, vline4,12 = 1-90.05em, hline1,10 = -0.1em, hline3,9 = -0.05em, cell3-45,7,9,11,13 = bg=customred, cell55 = bg=customgreen, cell57,9,11,13 = bg=customred, cell65,7,9,11,13 = bg=customred, cell75,9,11 = bg=customgreen, cell77,13 = bg=customred, cell85,9,11,13 = bg=customgreen, cell87 = bg=customred, cell95,7,9,11,13 = bg=customgreen, Case Classic Dynamic Clean-label Clean-lmage Average
+Case Classic Dynamic Clean-label Clean-lmage Average
 $\mathcal{L}_{\text{un}}$ $\mathcal{L}_{\text{re}}$ $\mathcal{L}_{\text{distill}}$ CA ASR CA ASR CA ASR CA ASR CA ASR
-✓ 83.6 97.5 90.2 82.0 91.7 98.5 87.1 97.3 88.1 93.8
+83.6 97.5 90.2 82.0 91.7 98.5 87.1 97.3 88.1 93.8
 93.2 91.5 93.6 62.2 93.6 94.4 91.2 97.2 92.9 86.3
-✓ 26.8 13.8 51.8 64.5 34.0 29.3 24.0 25.7 34.1 33.3
-✓ ✓ 93.1 48.3 93.5 34.5 93.6 93.8 92.0 37.2 93.0 53.5
-✓ ✓ 9.3 1.1 36.8 58.9 13.5 14.3 9.7 14.3 17.3 22.1
-✓ ✓ 92.8 4.1 93.6 24.5 92.9 11.8 92.2 13.7 92.9 13.5
-✓ ✓ ✓ 93.0 0.0 93.5 0.3 93.4 0.0 92.4 0.2 93.1 0.1
+26.8 13.8 51.8 64.5 34.0 29.3 24.0 25.7 34.1 33.3
+93.1 48.3 93.5 34.5 93.6 93.8 92.0 37.2 93.0 53.5
+9.3 1.1 36.8 58.9 13.5 14.3 9.7 14.3 17.3 22.1
+92.8 4.1 93.6 24.5 92.9 11.8 92.2 13.7 92.9 13.5
+93.0 0.0 93.5 0.3 93.4 0.0 92.4 0.2 93.1 0.1
 Table 6. CGD using other language-image pretraining models except CLIP.
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[260]Q[138]Q[138]Q[160]Q[160], cells = c, cell12 = c=2, cell14 = c=2, hline2 = 2,4l, hline2 = 3,5r, hline1,8 = -0.1em, hline3,7 = -0.05em, Model→ SigLIP ImageBind
+Model→ SigLIP ImageBind
 Attack↓ CA ASR CA ASR
@@
 Models
-Although we mainly use CLIP (Radford et al., 2021) to illustrate CGD’s effectiveness, CGD can use other language-image pretraining models to replace CLIP.
-As shown in Table 6, we consider SigLIP (Zhai et al., 2023) and ImageBind (Girdhar et al., 2023) as two alternative models.
-The result suggests that both achieve low ASR ($\leq 5\%$) while maintaining good CA ($\geq 90\%$).
-Figure 6. CGD’s defense performance on low-accuracy CLIP models. CGD successfully defends against all attacks (ASR $\leq 5\%$) even when CLIP’s zero-shot accuracy is as low as 10%.
+[…]
+[…]
 5.4. Security Guarantee Analysis
@@
 On these datasets, we tested three representative attacks (Blend, BPP, CTRL) but excluded clean-image backdoors due to either low ASR on large datasets (FLIP) or high computational cost (GCB).
-As shown in Table 7, CGD successfully defends all attacks except on SVHN for BPP.
+As shown in Table 5.4.1, CGD successfully defends all attacks except on SVHN for BPP.
 This failure is because CLIP’s zero-shot accuracy on SVHN is only 13.4%, making it thoroughly ineffective for a 10-class prediction dataset.
@@
 CGD fails only when CLIP is thoroughly ineffective (e.g., on SVHN) on dynamic backdoor attack.
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[221]Q[92]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83], cells = c, cell11 = r=2, cell12 = r=3, cell13 = c=2,r=2, cell15 = c=6, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell43 = bg=customblue17, cell44 = bg=customblue0, cell53 = bg=customblue16, cell54 = bg=customblue17, cell63 = bg=customblue13, cell64 = bg=customblue0, cell73 = bg=customblue11, cell74 = bg=customblue7, cell83 = bg=customblue11, cell84 = bg=customblue5, cell93 = bg=customblue10, cell94 = bg=customblue0, cell103 = bg=customblue8, cell104 = bg=customblue9, cell113 = bg=customblue8, cell114 = bg=customblue9, cell123 = bg=customblue7, cell124 = bg=customblue7, cell133 = bg=customblue7, cell134 = bg=customblue5, cell143 = bg=customblue7, cell144 = bg=customblue5, cell153 = bg=customblue6, cell154 = bg=customblue2, cell163 = bg=customblue3, cell164 = bg=customblue1, cell173 = bg=customblue2, cell174 = bg=customblue1, cell183 = bg=customblue2, cell184 = bg=customblue5, cell193 = bg=customblue2, cell194 = bg=customblue1, cell203 = bg=customblue0, cell204 = bg=customblue0, cell4-206 = bg = customgreen, cell5-208 = bg = customgreen, cell48 = bg = customred, cell4-2010 = bg = customgreen, hline3 = 3,5,7,9,11l, hline3 = 4,6,8,10,12r, hline2 = 5-11, hline1,21 = -0.1em, hline4 = -0.05em, Task→ Data
-Size
-Classify
-Accuracy
-CGD (Ours)
+Task→ Data
+Size Classify
+Accuracy CGD (Ours)
 Blend BPP CTRL
@@
 5.5.3. Results.
-Table 9 demonstrates that our CGD defense effectively mitigates both adaptive attacks.
-Against Adaptive-SIG, CGD reduces ASRs to $\leq 2.4\%$.
-As shown in Fig.
-7, while some triggered samples evade detection during the splitting stage, they retain minimal trigger information, with strongly triggered samples all identified and unlearned.
-For the FMB attack, ASR is also reduced to $\leq 1.2\%$ across both one-to-one and all-to-one settings.
-This occurs because, although the CLIP model may confuse mixed-feature samples, it only misclassifies them within the two mixed classes, not the target (third) class.
-Consequently, the poisoned target label also yields a high cross-entropy score, making the attack ineffective against our defense.
+[…]
 Table 9. Potential adaptive attacks against our CGD.
-{tblr}
-width = 0.8rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[250]Q[250]Q[110]Q[110]Q[110]Q[110], cells = c, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell21 = c=2, cell31 = r=3, cell61 = r=2, hline1,8 = -0.1em, hline2 = 3,5l, hline2 = 4,6r, hline3,6 = -0.05em, Defense → No Defense CGD (Ours)
+Defense → No Defense CGD (Ours)
 Adaptive Attacks ↓ CA ASR CA ASR
@@
 A.1. Detection of Label-Poisoned Samples with CLIP
-Consider a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, where a fraction $pr$ ($10<pr<$ 1) of samples are poisoned with incorrect labels.
-The clean subset is $C$ ($|C|=(1-pr)N$), and the poisoned subset is $P$ ($|P|=prN$).
-For each sample, CLIP computes a cross-entropy score $\mathcal{S}_{i}^{\text{CLIP}}=-\log p(y_{i}\mid x_{i})$, where $p(y_{i}\mid x_{i})$ is the predicted probability of the assigned label $y_{i}$.
-We rank samples by $\mathcal{S}_{i}^{\text{CLIP}}$ in descending order and flag the top $k=prN$ as potentially poisoned.
+[…]
 A.1.1. Calibration Assumptions and Score Behavior
@@
 Table 10. Attack performance under different poison rates.
-{tblr}
-width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.10em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Poison Rate 1% 10%
+Poison Rate 1% 10%
 Attack $\downarrow$ No Defense CGD (ours) No Defense CGD (ours)
@@
 Table 11. Attack performance under different architectures.
-{tblr}
-width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.08em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Architecture VGG16 MobileNet V2
+Architecture VGG16 MobileNet V2
 Attack $\downarrow$ No Defense CGD (ours) No Defense CGD (ours)
@@
 Table 12. Ablation study for poisoned dataset splitting with entropy from different models.
-{tblr}
-width = 1.0rowsep = 0.5pt, colsep = 0.5pt, colspec = Q[150]Q[60]Q[60]Q[60]Q[60]Q[80]Q[80]Q[80]Q[80]Q[60]Q[60], cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell3-53 = bg=customgreen, cell3,55 = bg=customgreen, cell45 = bg=customred, cell4,57 = bg=customgreen, cell37 = bg=customred, cell3,59 = bg=customgreen, cell49 = bg=customred, cell3,511 = bg=customgreen, cell411 = bg=customred, vline2,10 = 1-50.05em, hline1,6 = -0.1em, hline3 = -0.05em, Category → Classic Dynamic Clean-label Clean-image Average
+Category → Classic Dynamic Clean-label Clean-image Average
 Method ↓ CA ASR CA ASR CA ASR CA ASR CA ASR
@@
 For the sole exception, SVHN (vanilla CLIP accuracy: 13.4%, near-random guess for 10-class dataset), switching to LAION-2B CLIP variant reduced the ASR to 2.6% under the BPP attack.
-In other domain-specific environments like medical imaging, CLIP models such as PubMed (Eslami et al., 2023) (62.7% accuracy on PCAM) achieved an ASR of 6.1%, as shown in Table 13.
+In other domain-specific environments like medical imaging, CLIP models such as PubMed (Eslami et al., 2023) (62.7% accuracy on PCAM) achieved an ASR of 6.1%, as shown in Table F.
 These results demonstrate that with advanced domain-specific CLIP variants (e.g.
@@
 Table 13. Performance of different CLIP pre-training weights on SVHN and PCAM.
-{tblr}
-width = 0.8rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[238]Q[254]Q[192]Q[113]Q[113], cells = c, cell12 = r=2, cell13 = r=2, cell14 = c=2, cell31 = r=3, cell61 = r=3, hline1,9 = -0.1em, hline3,6 = -0.05em, Task→ CLIP Weight CLIP Acc.
-BPP Attack
+Task→ CLIP Weight CLIP Acc. BPP Attack
 Dataset↓ CA ASR
-SVHN
 (10 classes) Vanilla 13.4 94.3 100
@@
 LAION-2B (Schuhmann et al., 2022) 38.8 95.6 2.6
-PCAM
 (2 classes) Vanilla 52.3 82.1 69.5
@@
 In this case, poisoned-data-based defenses can get better defense performance than clean-data-based defenses with the help of “SC” (Split Clean data).
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,12-1412 = bg = customgreen, cell5,7,10,11,1512 = bg = customred, cell4,6-1514 = bg = customgreen, cell514 = bg = customred, cell4,6-9,11-13,1516 = bg = customgreen, cell5,10,1416 = bg = customred, cell4-13,1518 = bg = customgreen, cell1418 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 100% Poisoned Training Data
+Requirement → 5% Clean Training Data 100% Poisoned Training Data
 Defense → FP (Liu et al., 2018) ANP (Wu and Wang, 2021) I-BAU (Zeng et al., 2021) MCR (Zhao et al., 2020) FP-SC ANP-SC I-BAU-SC MCR-SC
 Attack ↓ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
-Classic
 Backdoor BadNets 92.3 2.1 85.6 0.0 86.2 0.8 92.1 2.4 93.3 1.0 85.4 0.0 90.1 1.5 91.2 1.0
 Blend 91.9 41.4 85.0 42.7 86.5 32.3 93.1 99.8 93.3 35.3 87.9 37.6 91.3 49.9 90.7 1.9
-Dynamic
 Backdoor WaNet 93.2 5.4 82.6 0.6 90.3 3.8 93.4 12.9 94.5 2.1 85.5 0.1 91.0 5.1 92.1 1.1
@@
 SSBA 92.0 14.5 87.3 0.1 90.0 1.5 92.4 64.2 92.9 7.7 86.1 0.0 91.2 4.3 90.5 1.0
-Clean-Label
 Backdoor CTRL 92.5 98.5 93.0 3.8 89.5 25.4 93.5 99.3 93.1 78.5 86.1 1.9 90.2 41.2 90.6 2.5
@@
 LC 92.1 34.7 87.6 51.1 88.9 5.3 93.6 100.0 93.1 19.3 85.5 16.7 92.1 1.6 90.4 1.4
-Clean-Image
 Backdoor FLIP 91.9 0.1 91.7 0.1 87.9 0.3 91.9 0.3 93.1 12.7 88.0 0.0 90.1 0.3 90.6 0.3
@@
 In this case, poisoned-data-based defenses can get comparable defense performance to clean-data-based defenses with the help of “SC” (Split Clean data).
-{tblr}
-width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,1312 = bg = customgreen, cell5,7,10-12,14,1512 = bg = customred, cell4,6-11,13-1514 = bg = customgreen, cell5,1214 = bg = customred, cell4,7-13,1516 = bg = customgreen, cell5,6,1416 = bg = customred, cell6-818 = bg = customgreen, cell4,5,9-1518 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 10% Poisoned Training Data
+Requirement → 5% Clean Training Data 10% Poisoned Training Data
 Defense → FP (Liu et al., 2018) ANP (Wu and Wang, 2021) I-BAU (Zeng et al., 2021) MCR (Zhao et al., 2020) FP-SC ANP-SC I-BAU-SC MCR-SC
 Attack ↓ CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
-Classic
 Backdoor BadNets 92.3 2.1 85.6 0.0 86.2 0.8 92.1 2.4 92.5 2.7 85.0 0.0 89.5 1.2 92.8 34.1
 Blend 91.9 41.4 85.0 42.7 86.5 32.3 93.1 99.8 92.5 37.6 88.4 42.5 89.0 59.4 93.5 99.3
-Dynamic
 Backdoor WaNet 93.2 5.4 82.6 0.6 90.3 3.8 93.4 12.9 89.6 2.1 82.2 0.1 91.8 22.1 93.1 0.6
@@
 SSBA 92.0 14.5 87.3 0.1 90.0 1.5 92.4 64.2 88.6 16.1 86.0 0.0 90.2 10.9 92.5 51.5
-Clean-Label
 Backdoor CTRL 92.5 98.5 93.0 3.8 89.5 25.4 93.5 99.3 91.1 95.3 93.2 2.3 90.5 6.5 93.6 99.6
@@
 LC 92.1 34.7 87.6 51.1 88.9 5.3 93.6 100.0 92.1 47.8 89.3 56.6 88.5 7.1 92.9 100.0
-Clean-Image
 Backdoor FLIP 91.9 0.1 91.7 0.1 87.9 0.3 91.9 0.3 92.1 1.9 90.5 0.0 88.2 3.4 92.5 76.7
@@
 In this section, we evaluate clean-data based defenses under three conditions: 5% clean training dataset, whole (100%) poisoned training dataset, and 10% poisoned training dataset.
-Table 14 presents the results of clean-data based defenses with 5% clean training data available.
+Table F presents the results of clean-data based defenses with 5% clean training data available.
 In this setting, the total number of successful defenses (defined as defenses achieving an Attack Success Rate (ASR) of 20% or lower, highlighted in green) is 23.
-Table 14 shows the results for clean-data based defenses when a clean subset is split from the whole (100%) poisoned training dataset.
+Table F shows the results for clean-data based defenses when a clean subset is split from the whole (100%) poisoned training dataset.
 With access to the full poisoned dataset, the defense performance improves significantly, with the number of successful defenses increasing to 35, marking a 50% improvement over the 5% clean data setting.
 This indicates that having access to the full poisoned dataset allows the defenses to operate more effectively, even in the absence of originally clean data.
-Table 15 provides results for clean-data based defenses with a 10% poisoned training dataset.
+Table F provides results for clean-data based defenses with a 10% poisoned training dataset.
 Even in this setting, where only a small portion of the dataset is clean, the total number of successful defenses increases from 23 (in the 5% clean data setting) to 25.